### v1

# Anticipating Gaming to Incentivize Improvement: Guiding Agents in (Fair) Strategic Classification

Abstract
As machine learning algorithms increasingly influence critical decision making in different application areas, understanding human strategic behavior in response to these systems becomes vital. We explore individuals’ choice between genuinely improving their qualifications (“improvement”) vs. attempting to deceive the algorithm by manipulating their features (“manipulation”) in response to an algorithmic decision system. We further investigate an algorithm designer’s ability to shape these strategic responses, and its fairness implications. Specifically, we formulate these interactions as a Stackelberg game, where a firm deploys a (fair) classifier, and individuals strategically respond. Our model incorporates both different costs and stochastic efficacy for manipulation and improvement. The analysis reveals different potential classes of agent responses, and characterizes optimal classifiers accordingly. Based on these, we highlight the impact of the firm’s anticipation of strategic behavior, identifying when and why a (fair) strategic policy can not only prevent manipulation, but also incentivize agents to opt for improvement.
footnotetext: Department of Integrated Systems Engineering, The Ohio State University, alhanouti.1@osu.edu. Department of Electrical and Computer Engineering, University of California, San Diego, parinaz@ucsd.edu.
Keywords: strategic classification, game theory, fair machine learning, manipulation, improvement.
1 Introduction
Machine learning (ML) algorithms have come to play a pivotal role in guiding decision making in many application areas, including banking, hiring, social media, and resource allocation. While the use of ML-driven systems can enhance efficiency, it can also drive the humans who are subject to algorithmic decisions to adjust their behavior accordingly. Examples include Uber drivers coordinating their behavior in response to its surge pricing algorithm [Möhlmann and Zalmanson, 2017], applicants selecting keywords and formatting to pass automated resume screening [Forbes, 2022], and Facebook users adjusting their posting and content interaction choices in response to the platforms’ curation algorithms [Eslami et al., 2016]. These can be viewed as strategic responses by rational human subjects in these systems, motivating a game-theoretical analysis of learning algorithms with human in the loop.
Earlier works on the study of strategic humans facing ML systems largely focused on scenarios where users can strategically alter only their observable data (e.g., students cheating to obtain better test scores, job applicants making formatting or wording changes to their CV, or loan applicants opening several new accounts to increase their credit scores) to receive a favorable decision (e.g., be accepted to a school, job opening, or loan); see, e.g., [Hu et al., 2019, Milli et al., 2019]. This strategic behavior is referred to as strategic manipulation, where agents change their features without changing their true qualification states. This can be interpreted as cheating the machine learning algorithm: such agents may appear to be more qualified, without being truly suitable for a favorable outcome.
However, strategic responses by human agents need not be limited to manipulation: some agents may choose to put in the effort to truly change their qualification state, instead of simply cheating the system; examples include students studying for (standardized) tests to attain high scores, or loan applicants working on their finances to improve their ability to repay a loan.
This type of strategic behavior, which changes the (observable) features as well as the (unobservable) true qualification state in an attempt to become qualified for a favorable outcome is referred to as strategic improvement; this aspect has been receiving increased attention in the subsequent works on strategic machine learning, e.g., [Haghtalab et al., 2020, Miller et al., 2020, Kleinberg and Raghavan, 2020, Efthymiou et al., 2025]. Generally, from the viewpoint of any institution, the preferred strategic behavior is for users to invest in improvement rather than to try to manipulate their system. Therefore, it is of interest to understand the incentives of strategic agents to opt for each type of response, its implications for the institution’s choice of a decision policy, and potential ways to shape users’ strategic behavior. In this paper, we explore such questions.
In particular, we consider a strategic classification problem, where strategic agents facing a binary classier can decide between exerting effort to either manipulate and only change their features, or improve by changing both their feature and their true qualification state, with the goal of being selected by the classifier. We allow these actions to have different costs, and also differ in the resulting change in the agents’ initial features, with the efficacy of each action being stochastic; as we detail in Section 2, these stochastic efficacies are one of the distinctions of our model from prior works that have accounted for both manipulation and improvement actions in strategic classification. Our goal is to understand the effects of these different costs and efficacies on the agents’ choice to opt for each decision, as well as understand how a firm can leverage this understanding to modify its algorithms to encourage improvements while discouraging manipulations.
To this end, we formulate this setting as a Stackelberg game where the firm first deploys a (fair) classifier, and then agents strategically respond to it. We first characterize the Stackelberg game equilibrium (Proposition 2) under the assumption that improvement is more costly than manipulation, but that it is more effective at advancing agents’ chances of receiving a favorable decision (we do the latter by assuming that the transition probabilities under manipulation first-order stochastically dominate those under improvement). We then show (Proposition 8) how the firm’s knowledge of the strategic behavior (or lack thereof) impacts its choice of a classifier, as well as agents’ strategic choices and outcomes across different qualification states. The main challenge in finding these optimal classifiers is that the regions in which agents opt for the improvement and manipulation actions are themselves functions of the firm’s decision variables, and these dependencies should be accounted for when evaluating the firm’s best-response. This is done by applying the Leibniz integral rule to the firm’s objective function (with adaptations when the firm’s optimal classifier selection problem is subject to fairness constraints), and then leveraging the relation between the threshold and the indifference points in agents’ strategic responses, which appear in the integration limits and their derivatives, to simplify the resulting expressions of the firm’s problem’s optimality conditions.
Building on the characterization of the firm’s optimal policies, we then provide additional insights on the impacts of anticipating agent’s strategic behavior, and the impacts of fairness interventions, on shaping the agents’ responses. Specifically, in Section 6.1, we first discuss how anticipating strategic behavior allows the firm to not only curb undesired manipulation (as also found in prior work) but can also incentivize all agents, whether initially qualified or unqualified, to opt for improvement decisions instead. Notably, we find that the firm will make its selection algorithm more “strict”, incentivizing improvement by unqualified agents (driving them to improve both their qualification states and observable features), while still leaving the manipulation option open to qualified agents (who do not have sufficiently high observable features to be selected otherwise, but whose acceptance benefits the firm). Together, these findings lead to our first key takeaway: anticipating agents’ strategic responses, which the firm does solely with the goal of optimizing its utility, has the added benefit of guiding agents’ behavior in a desirable direction.
Then, in Section 6.2, we investigate the implications of adopting algorithmic fairness interventions, when both improvement and manipulation actions are available to the participating agents. We argue that in its attempt to attain fair outcomes, a firm may decrease improvement incentives for agents from a disadvantaged population; this can be viewed as a negative consequence of fairness interventions on shaping agents’ strategic responses, and emerges even if the differences between demographic groups are only historical (in qualification rates and feature distributions) and not current (in access to strategic resources). As we elaborate through numerical experiments in Section 7, this negative by-product is more severe if the firm is non-strategic (i.e., unaware of agents’ strategic responses). In comparison, a strategic firm recognizes that fairness goals can be met by a combination of adjusting decision thresholds and incentivizing improvement actions. In fact, we observe that fairness desiderata can be achieved with very small changes to the decision thresholds when accounting for agents’ strategic responses, therefore leading to only minor changes in agents’ improvement incentives, and far less loss in utility for the firm. Together, these findings lead to our second key takeaway: existing literature on fairness interventions (e.g., Liu et al. [2018], Hu et al. [2019], Estornell et al. [2023], Keswani and Celis [2023]) often find that fairness is improved in two main ways: inclusiveness (flipping disadvantaged group predictions to positive) or selectivity (flipping advantaged group predictions to negative); we identify a third lever: anticipating strategic responses, which increases improvement incentives, and in turn aids in improving fairness “for free”.
To summarize, our findings offer two new insights for algorithm designers, as well as for managers contemplating the deployment of algorithmic decision rules within their organizations, in the presence of strategic users. First, anticipating the users’ strategic responses to algorithmic systems (which are increasingly observed, as noted in the examples in the beginning of this section) can not only help protect the planned operating points of these systems, but also will ultimately help enhance the system by increasing the number of agents (workers, employees, students, etc.) who select the preferred improvement actions over undesired manipulation actions, in turn leading to an increase in users’ qualification/success levels. Second, as developing fair and responsible algorithmic decision systems is becoming of increased interest due to reputation effects and applicable laws, driving agents’ improvement actions can be viewed as an additional lever in meeting these fairness goals, allowing firms to attain fairness desiderata without having to lower accuracy significantly.
In the remainder of the paper, we begin with a review of related work in Section 2. We then introduce our proposed strategic classification model in Section 3. In Section 4, we characterize agents’ strategic responses under this model, and the resulting post-strategic population statistics, given a fixed classifier. We then characterize the firm’s optimal choice of classifier in Section 5, when the firm is aware or unaware of the impacts of its choice on agents’ responses, as well as with and without fairness interventions. We theoretically analyze the effects of anticipating strategic behavior and the inclusion of fairness constraints in Section 6, followed by numerical experiments and illustrations supporting our analytical findings in Section 7. We conclude with potential directions for future work in Section 8.
2 Related Work
“Strategic” machine learning, and in particular, strategic classification and regression problems, wherein ML/AI algorithm designer(s) and/or users (rationally) adjust their decisions in response to others’, has become a subject of growing interest. A majority of the research in this area studies settings where humans adjust their behavior with the goal of receiving more favorable algorithmic decisions, and falls within one of three main categories: (i) agents strategically manipulating observable data (features) to secure favorable outcomes without changing their true qualifications [Cohen et al., 2025, Levanon and Rosenfeld, 2022, Sundaram et al., 2023, Cohen et al., 2024, Lechner et al., 2023, Ahmadi et al., 2023, Shao et al., 2023, Levanon and Rosenfeld, 2021, Zhang et al., 2022, Ahmadi et al., 2021, Ebrahimi et al., 2024] and (ii) agents choosing between manipulation and improvement actions, where improvement changes their true qualification [Jin et al., 2023, Wang et al., 2023, Horowitz and Rosenfeld, 2023, Haghtalab et al., 2020, Jin et al., 2022, Miller et al., 2020, Zhang et al., 2020b, Ahmadi et al., 2022, Kleinberg and Raghavan, 2020, Bechavod et al., 2021, Shavit et al., 2020, Harris et al., 2021, Alon et al., 2020, Bechavod et al., 2022, Xie et al., 2024, Efthymiou et al., 2025] . Our work falls within the third category, and differs from the prior literature mainly in our model of both distinct costs as well as different stochastic efficacy for these strategic actions, as well as our exploration of fairness implications of the availability of these actions. We elaborate on these differences below. Before proceeding, we also note that our work (and the closely related literature) touches upon one set of strategic actions; other strategic actions available to firms/agents include strategic interactions between multiple firms [Narang et al., 2022] or agents themselves [Eilat et al., 2022], agents strategically hiding data from the classifier instead of distorting data [Zhang et al., 2020a], or firms adopting randomized classifiers [Braverman and Garg, 2020, Geary and Gouk, 2025].
Among the works that consider manipulation and improvement actions as those available to agents, Shavit et al. [2020] and Bechavod et al. [2021] model different actions available to agents by defining a set of features that may or not affect the true qualification when they change, and propose a conversion matrix to represent the effect of the actions accordingly. Similarly, Horowitz and Rosenfeld [2023],Wang et al. [2023], and Miller et al. [2020] use causal modeling to differentiate actions available to agents that affect the true qualification versus those that only alter predictions. Additionally, Jin et al. [2023], like our work, distinguish between manipulation and improvement actions, but introduce a causal model where an agent’s endowed pre-response attribute causally influences both observable features and the true qualification, with effects captured by a projection matrix.
In contrast, we consider distinct manipulation and improvement actions, with their effects stochastically following different probability distributions. Emelianov et al. [2022], similar to our work, consider the effect of the actions to be probabilistic, but following a normal distribution around the selected effort; in contrast, our proposed model of stochastic efficacy of actions does not specify the boost distributions.
Lastly, our work is also closely related to those that explore the fairness implications of the availability of strategic actions to users. The existing work in this area (e.g., [Milli et al., 2019, Hu et al., 2019, Zhang et al., 2022, Estornell et al., 2023]) study the fairness implications when agents can only manipulate their features. The impacts of the availability of both types of strategic behavior, to our knowledge, remains unexplored. Our work addresses this gap, providing insights that firms can leverage to adapt their algorithms in a way that attains fairness desiderata, in part, by incentivizing improvement decisions.
An early version of this work appeared in [Alhanouti and Naghizadeh, 2024], where we first proposed our model of strategic classification with manipulation and improvement decisions, and provided a preliminary analysis of the impacts of anticipating agents’ strategic behavior. The current paper provides an expanded analysis of anticipating agents’ responses (Proposition 8), an analysis of fairness implications (Corollary 9, and related discussions), and detailed numerical verification and illustrations (Section 7).
3 Problem Setting and Preliminaries
We consider a Stackelberg game where the firm (algorithm designer) first announces a classifier, following which the agents respond strategically. In this section, we detail the agents’ and the firm’s characteristics, their actions, and their utilities. Our model is similar to, and extends, those of prior work on strategic classification Zhang et al. [2020b, 2022]. The notation is summarized in Table 1.
Table 1: Summary of main notation.
Notation Description
$s\in\{a,b\}$ Demographic group membership.
$n_{s}\in[0,1]$ Fraction of agents in group $s$.
$x\in\mathbb{R}_{\geq 0}$ Agents’ observable feature pre-strategic behavior.
$y\in\{0,1\}$ Agents’ unobservable, true qualification state pre-strategic behavior.
$\alpha_{s}\in[0,1]$ The qualification rate in group $s$ pre-strategic behaviour.
$G^{y}_{s}(x)$ Feature distribution (pdf) of qualification $y$, group $s$, agents, pre-strategic behavior.
$\hat{x},\hat{y}$, $\hat{\alpha}_{s},\hat{G}_{s}^{y}(x)$ Agents’ feature, qualification state, group qualification rate, and feature distribution, post-strategic behavior.
$w\in\{M,I,N\}$ Agents’ strategic actions: manipulation ($M$), improvement ($I$), and doing nothing ($N$).
$C_{w,s}$ Cost of strategic action $w$ to an agent from group $s$.
$\tau^{y}_{w,s}(b)$, $\mathbb{T}^{y}_{w,s}(b)$ The pdf and CDF of the distribution of the boost $b$ from action $w$ for an agent from group $s$ with true qualification $y$.
[…]
$\theta_{s}$ The firm’s decision threshold for group $s$.
$u_{+}/u_{-}$ The firm’s benefit/cost from true/false positives.
$\mathbb{I}^{y}_{s}/\mathbb{M}^{y}_{s}/\mathbb{N}^{y}_{s}$ Set of agents from group $s$ with qualification $y$ who opt for improvement/manipulation/doing nothing.
The agents.
Consider a population of agents with two types of observable attributes: sensitive features that divide the population into two demographic groups $S\in\{a,b\}$ (e.g., gender dividing population into Male and Female, or race dividing the population into White and Non-White agents), and an (observable) feature/score $x\in\mathbb{R}_{\geq 0}$ (e.g., a credit score, an exam score). Let $n_{s}$ be the fraction of agents in group $s$.
Each agent has a true (hidden) label or qualification state, denoted $Y\in\{0,1\}$, with $Y=1$ denoting a qualified agent (e.g., one who can pay off a loan if granted) and $Y=0$ denoting an unqualified agent. Let $\alpha_{s}:=\mathbb{P}(Y=1|S=s)$ denote the qualification rate in group $s$. In addition, let $G^{y}_{s}(x):=\mathbb{P}(X=x|Y=y,S=s)$ denote the probability density function (pdf) of the distribution of the features for individuals with qualification state $y$ from group $s$. We make the following assumptions on the feature distributions.
Assumption 1.
The feature distributions $G^{y}_{s}(x)$ are continuous, and satisfy the strict monotone likelihood ratio property: $\frac{G^{1}_{s}(x)}{G^{0}_{s}(x)}$ is strictly increasing in $x\in\mathbb{R}_{\geq 0}$.
In words, this entails that as an agent’s feature $x$ increases, the likelihood that the agent is qualified increases.
The firm.
A firm makes decisions on agents based on their observable features $x$ and their group memberships $s$. The decision is binary, denoted $d\in\{0,1\}$, where $d=1$ represents acceptance and $d=0$ represents rejection. This decision is determined by a group-dependent binary classifier $\pi_{s}(x)=\mathbb{P}(D=1|X=x,S=s)$. We assume that this is a threshold policy, such that an agent is accepted if and only if ${x}\geq\theta_{s}$. (Zhang et al. [2020b] show that such threshold policies are optimal under mild assumptions.)
Agents’ strategic actions.
After the threshold policy $\theta_{s}$ is announced by the firm, agents have the option to behave strategically to improve their chances of receiving favorable outcomes. This is done by choosing one of the strategic actions $w\in\{M,I,N\}$, with $M$ denoting manipulation, $I$ denoting improvement, and $N$ denoting doing nothing.
Taking these actions may impact the agents’ features and/or true labels. We use $X$ and $Y$ to denote the pre-strategic feature and true label (before an action $w$ is taken), and $\hat{X}$ and $\hat{Y}$ to denote the post-strategic ones. In particular, both qualified and unqualified agents ($y=1$ and $y=0$) can opt to manipulate ($w=M$) by changing their feature $x$ to some new feature $\hat{x}$, without changing their true qualification state ($\hat{y}=y$); e.g., this could represent cheating on an exam, which increases the score without affecting true learning. Alternatively, both types of agents can choose to invest in improving themselves ($w=I$), which leads to a change in both their feature from $x$ to some $\hat{x}$, as well as their true label to $\hat{y}=1$; e.g., this could represent studying for an exam, which improves both the score and true learning. Lastly, agents who opt to do nothing ($w=N$) maintain their feature and label ($\hat{x}=x$ and $\hat{y}=y$). The firm will only observe the agents’ post-strategic features $\hat{x}$ when making its decision. We let $\hat{\alpha}_{s}$ and $\hat{G}_{s}^{y}(x)$ denote the post-strategic population statistics after these changes have happened as a result of agents’ selected strategic actions. (We will formally characterize these updated statistics in Section 4.2).
In addition to differing in their impacts on changes in features and/or true labels, the actions differ in two aspects: their cost and their efficacy. First, each action requires exerting effort and comes at a certain group-dependent (constant) cost ${C_{w,s}}\in[0,1)$, $w\in\{M,I,N\},s\in\{a,b\}$. Such difference in costs across actions and groups is a common assumption in prior works on strategic classification (e.g., Zhang et al. [2022], Liu et al. [2020], Jin et al. [2022]). The element that is new to our model is that we assume all actions lead to a (weak) increase in the feature (i.e., we assume that ${\hat{x}\geq x}$), but that this increase in feature is stochastic and different across actions and groups. Formally, we assume the probability that the feature $x$ of a qualified/unqualified individual from a group $s$ increases by $b_{w}$ after opting for action $w$ is distributed according to a boost distribution with pdf ${\mathbb{\tau}^{y}_{w,s}(b)}:=\mathbb{P}(\hat{X}=x+b|X=x,Y=y,W=w,S=s)$. These boost distributions determine the efficacy of the action. As an example, they could capture that both cheating and studying can increase an agent’s exam score, but that the efficacy of these actions in improving scores is different. Let $\{\underline{b}^{y}_{w,s},\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu^{y}_{w,s}\}$ denote the minimum and maximum boost under action $w$, and $\mathbb{T}^{y}_{w,s}(b)$ denote the CDF of the boost function.
We let $C_{N,s}=0$ and $\mathbb{\tau}^{y}_{N,s}(0)=1$, meaning that the “do nothing” action has zero cost and no impact on changing the agent feature. The two remaining actions, $M$ and $I$, can differ in cost and efficacy. We conduct our analysis under the following assumption.
Assumption 2.
Improvement is more costly than manipulation (i.e., $C_{I,s}\geq C_{M,s}$), and the improvement boost distribution first-order stochastically dominates (FOSD) that of manipulation (i.e., $\mathbb{T}^{y}_{M,s}(b)\geq\mathbb{T}^{y}_{I,s}(b),\forall b$).
In words, an action first-order stochastically dominating another means that it “gets the agent further”, in that it has a higher probability of increasing the agent’s feature from $x$ to a feature greater or equal to $\hat{x}=x+b$. Assumption 2 gives rise to the conflict between the two actions: manipulation is cheaper, while improvement is more effective in advancing the agent. (We note that there are three scenarios beyond Assumption 2: one can be recovered by interchanging the indexes $M$ and $I$ in our results. The remaining two cases are less interesting, as one of the actions would be more beneficial in both cost and efficacy.)
Agents’ utility.
In general, each agent’s strategic choice among the three available actions (manipulation, improvement, or doing nothing) depends on its budget $B$, the cost of effort $C_{w,s}$, the efficacy of the selected action $\mathbb{\tau}^{y}_{w,s}$, and the firm’s deployed policy $\theta_{s}$. For simplicity, we assume the same budget $B$ for all agents, and assume this budget is sufficiently high so that all agents can afford any action. Then, the choice among the actions depends on the relative benefit vs. the cost of each action.
Formally, an agent chooses to incur the cost $C_{w,s}$ of action $w$ if and only if it (sufficiently) increases the probability that the agent is accepted by the firm. For an agent from group $s$ with feature $x$ and true label $y$, the benefit from strategic action $w$ is the increase in the probability of being accepted, given by
$\mathcal{B}_{w,s}(x,y)$ $:=\mathbb{P}(D=1|X=x,Y=y,W=w,S=s)-\mathbb{P}(D=1|X=x,Y=y,W=N,S=s).$ (1)
Note that the efficacy of the selected action $\mathbb{\tau}^{y}_{w,s}$, and the firm’s deployed policy $\theta_{s}$, will affect this benefit. The agent’s overall utility is given by $u_{s}(x,y,w):=\mathcal{B}_{w,s}(x,y)-C_{w,s}$.
The firm’s utility.
The firm’s goal is to find an optimal policy that maximizes its expected utility by correctly classifying the agents. The firm receives benefit $u_{+}$ from accepting qualified individuals, and incurs penalty $u_{-}$ from accepting unqualified individuals. The firm’s goal may alternatively be finding the fair optimal policy by also imposing a fairness constraint $\mathcal{C}^{f}$ on its decision problem.
Formally, let $U(\theta_{a},\theta_{b})$ represent the firm’s total utility given the decision thresholds $\theta_{a}$ and $\theta_{b}$ for each group. Then, the firm’s expected utility is:
$\mathbb{E}[U(\theta_{a},\theta_{b})]$ $=\sum_{s\in\{a,b\}}n_{s}[\mathbb{P}(D=1,Y=1|S=s)u_{+}-\mathbb{P}(D=1,Y=0|S=s)u_{-}]$
$=\sum_{s\in\{a,b\}}n_{s}\int_{\theta_{s}}^{\infty}[G^{1}_{s}(x)\alpha_{s}u_{+}-G^{0}_{s}(x)(1-\alpha_{s})u_{-}]\mathrm{d}x~{}.$ (2)
The firm’s (fair) optimization problem is to choose the decision thresholds $\theta_{a}$ and $\theta_{b}$ as follows:
$\max_{\theta_{a},\theta_{b}}\quad\mathbb{E}[U(\theta_{a},\theta_{b})]~{},\qquad\textbf{s.t.}\quad\mathcal{C}^{f}_{a}(\theta_{a})=\mathcal{C}^{f}_{b}(\theta_{b}).$ (3)
The fairness constraint here, donated by $f$, will be considered to be one of the two more commonly studied notions of fairness Mehrabi et al. [2021]:
  • Equality of Opportunity (EOP): $\mathcal{C}^{EOP}_{s}=\int_{\theta_{s}}^{\infty}{G}_{s}^{1}(x)\mathrm{d}x$. This will equalize the true positive rates between groups $a$ and $b$.
  • Demographic Parity (DP): $\mathcal{C}^{DP}_{s}=\int_{\theta_{s}}^{\infty}({G}_{s}^{1}(x)\alpha_{s}+{G}_{s}^{0}(x)(1-\alpha_{s}))\mathrm{d}x$. This will equalize selection (acceptance) rates between groups $a$ and $b$.
When agents are strategic and this is known to the firm, the agents’ statistics in (2) will be replaced by the post-strategic values $\hat{\alpha}_{s}$ and $\hat{G}_{s}^{y}(x)$, as characterized shortly in Section 4.
4 Agents’ Strategic Behavior
We begin by characterizing agents’ best responses to a firm’s given classifier. Consider an agent from group $s$ with feature $x$ and true label $y$, facing decision threshold $\theta_{s}$. We will show that for any given $y$ and $\theta_{s}$, the feature space $x$ can be partitioned into disjoint regions determining the agents’ best-response, with the boundaries of these regions determined by the points where agents become indifferent between pairs of actions.
Specifically, we first define a set of indifference features based on the cost-efficacy trade-off of the actions available to the agents.
Definition 1.
Given efficacy CDFs $\mathbb{T}^{y}_{w,s}$, with inverse CDFs $(\mathbb{T}^{y}_{w,s})^{-1}$, and costs $C_{w,s}$, define:
  • Opt in features: $\mathbf{o}^{y}_{w,s}=\max\{0,\theta_{s}-(\mathbb{T}^{y}_{w,s})^{-1}(1-C_{w,s})\}$, for $w\in\{M,I\}$.
  • Flip decision feature: $\mathbf{f}^{y}_{s}\in[\theta-\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu_{M},\theta-\underline{b}_{I}]$ satisfying $\mathbb{T}^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})-\mathbb{T}^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})=C_{I,s}-C_{M,s}$, if a solution exists; zero otherwise.
  • Risk taker feature: $\mathbf{r}^{y}_{s}:=\max\{0,\theta_{s}-(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})\}$.
Intuitively, and as shown formally in the proof of Proposition 2, these features can be interpreted as follows: The opt in features determine the first feature at which the agents benefit from opting for an action $M$ or $I$ as opposed to doing nothing $N$; the flip decision feature is the feature at which the cost-efficacy trade-off of the $M$ and $I$ actions are equalized so that agents choices will flip between these decisions at this feature; and the risk taker feature is the first feature at which an agent opts for an uncertain shot at getting a favorable classification by choosing $M$ over the certain admission attainable from $I$, given $M$’s relatively lower cost.
Using the above, we characterize the agents’ optimal best responses to a given decision threshold $\theta_{s}$.
Proposition 2.
If $\mathbf{f}^{y}_{s}$ is unique, the agents’ optimal response $w^{*}_{s}(x,y)=\arg\max_{w\in\{M,I,N\}}u_{s}(x,y,w)$ to a given threshold $\theta_{s}$ will be one of the three types outlined in Table 2.
Table 2: Agents’ best responses (Proposition 2).
Type Condition Range : $w_{s}^{*}(x,y)$
Type 1 $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$ $[0,\mathbf{o}^{y}_{I,s}):N$, $[\mathbf{o}^{y}_{I,s},\mathbf{r}^{y}_{s}):I$, $[\mathbf{r}^{y}_{s},\theta_{s}):M$, $[\theta_{s},\infty):N$
Type 2 $\mathbf{o}^{y}_{M,s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{f}^{y}_{s}$ $[0,\mathbf{o}^{y}_{M,s}):N$, $[\mathbf{o}^{y}_{M,s},\mathbf{f}^{y}_{s}):M$, $[\mathbf{f}^{y}_{s},\mathbf{r}^{y}_{s}):I$, $[\mathbf{r}^{y}_{s},\theta_{s}):M$, $[\theta_{s},\infty):N$
Type 3 $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{o}^{y}_{I,s}$ $[0,\mathbf{o}^{y}_{M,s}):N$, $[\mathbf{o}^{y}_{M,s},\theta_{s}):M$, $[\theta_{s},\infty):N$
    Or $\mathbf{o}^{y}_{I,s}\leq\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$.
  • $[0,\mathbf{o}^{y}_{I,s}):N$, $[\mathbf{o}^{y}_{I,s},\mathbf{f}^{y}_{s}):I$, $[\mathbf{f}^{y}_{s},\theta_{s}):M$, $[\theta_{s},\infty):N$.
A detailed proof is given in Appendix A.1.1. These three types of best-responses are illustrated in Figure 1. In particular, note that in all types of equilibrium, agents who are close to the decision threshold opt to be risk takers, choosing uncertain but cheap manipulation over certain but costly improvement. Interestingly, we posit that this may present a parallel with gaming behavior observed in education settings: students committing academic dishonesty typically have higher GPAs [OEDB, n.d].
(a) Type 1.
(b) Type 2.
(c) Type 3.
Figure 1: Agent best-responses identified in Proposition 2.
We also note that the type of equilibrium is determined solely by the relative cost and efficacy of the manipulation and improvement actions; the choice of the classifier $\theta_{s}$ changes the indifference points where agents opt for each action, but not the type of equilibrium. We provide additional intuition for this below.
4.1 Illustration: uniform boost distributions
To further illustrate the intuition behind the types of best-responses identified in Proposition 2, we consider uniform boost distributions $\mathbb{\tau}^{y}_{w,s}$. It is straightforward to verify that the three possible equilibria of Proposition 2 can be obtained by varying improvement cost $C_{I}$ relative to manipulation cost $C_{M}$. Specifically, consider $\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu_{M}^{y}-\underline{b}_{M}^{y}\geq\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu_{I}^{y}-\underline{b}_{I}^{y}$. Then:
[…]
  • High improvement cost: If $C_{M}+\frac{\underline{b}_{I}^{y}-\underline{b}_{M}^{y}}{\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu_{M}^{y}-\underline{b}_{M}^{y}}\leq C_{I}$, the best-response is of Type 3 in Proposition 2. In this case, the improvement cost is significantly higher than the manipulation cost, leading all agents to pick manipulation over improvement once profitable.
4.2 Post-strategic population statistics
We next characterize the post-strategic population statistics. In particular, as noted earlier, the strategic actions $w\in\{M,I\}$ change the agent’s true qualification and/or feature. Denote the post-strategic qualification rates by $\hat{\alpha}_{s}=\mathbb{P}(\hat{Y}=1|S=s)$, and the post-strategic feature distribution by ${\hat{G}^{y}_{s}(x)}:=\mathbb{P}(\hat{X}={x}|\hat{Y}={y},S=s)$. Let $\mathbb{I}^{y}_{s,(i)}:=\{x|Y=y,S=s,w=I\}$ be the set of agents from group $s$ with label $y$ who opt for improvement under a given threshold policy $\theta_{s}$, in a type (i) best response (as identified in Proposition 2). Define $\mathbb{M}^{y}_{s,(i)}$ and $\mathbb{N}^{y}_{s,(i)}$ similarly for the set of agents who opt for manipulation and for doing nothing under equilibrium type (i).
The following lemma presents the expressions for the post-strategic population statistics in terms of the pre-strategic population statistics and the best-response regions.
Lemma 3.
The post-strategic qualification rate is given by
$\hat{\alpha}_{s}=\alpha_{s}+(1-\alpha_{s})\int_{x\in\mathbb{I}^{0}_{s,(i)}}G^{0}_{s}(x)\mathrm{d}x$ (4)
The post-strategic feature distributions are given by:
$\hat{G}^{0}_{s}(x)$ $=\frac{1-\alpha_{s}}{1-\hat{\alpha}_{s}}\Big{(}\mathbbm{1}(x\in\mathbb{N}^{0}_{s,(i)})G^{0}_{s}(x)+(G^{0}_{s}*\tau^{0}_{M,s})(x)\Big{)},$ (5)
$\hat{G}^{1}_{s}(x)$ $=\frac{\alpha_{s}}{\hat{\alpha}_{s}}\Big{(}\mathbbm{1}(x\in\mathbb{N}^{1}_{s,(i)})G^{1}_{s}(x)+(G^{1}_{s}*\tau^{1}_{M,s})(x)+(G^{1}_{s}*\tau^{1}_{I,s})(x)\Big{)}+\frac{1-\alpha_{s}}{\hat{\alpha}_{s}}(G^{0}_{s}*\tau^{0}_{I,s})(x),$ (6)
where $(G^{y}_{s}*\tau^{y}_{w,s})(x):=\int_{z\in\mathbb{W}^{y}_{s,(i)}}G^{y}_{s}(z)\tau^{y}_{w,s}(x-z)\mathrm{d}z$ is the convolution of the feature distribution and boost function restricted to the action $w$ region.
A detailed proof appears in Appendix A.1.2. Intuitively, (4) captures the increase in qualification rates due to improvement actions taken by previously unqualified individuals (those in $\mathbb{I}^{0}_{s}$). The new feature distribution of unqualified individuals in (5) consists of unqualified individuals who do nothing, and unqualified individuals who choose manipulation actions (with their features shifted by the boost distribution). The new feature distribution of qualified individuals in (6), on the other hand, consists of all qualified individuals (with appropriate feature shifts for those who opt for manipulation or improvement), as well as previously unqualified individuals who are now qualified due to opting for improvement actions.
5 The Firm’s Optimal Policy
Using the findings of Section 4, we can now determine the firm’s optimal choice of decision thresholds.
We start with a firm who does not account for agents’ strategic behavior and does not implement any fairness constraints; we refer to this as the unfair non-strategic firm. The lemma below characterizes such firm’s optimal decision thresholds as a function of the population statistics.
Lemma 4.
The unfair non-strategic firm’s optimal decision thresholds $\theta_{s}^{U}$ satisfies $\frac{G^{1}_{s}(\theta_{s}^{U})}{G^{0}_{s}(\theta_{s}^{U})}=\frac{u_{-}(1-\alpha_{s})}{u_{+}\alpha_{s}}$.
This is similar to results obtained in [Zhang et al., 2022, Liao and Naghizadeh, 2022]. The proof is provided in Appendix A.2.1 for completeness.
If the firm is cognizant of the agents’ strategic behavior, the unfair strategic firm’s optimal thresholds $\hat{\theta}_{s}^{U}$ can be obtained by finding the utility maximizer when the firm’s utility (2) is evaluated on the population post-strategic statistics characterized in Lemma 3. Specifically, the strategic firm’s utility can be written as:
$\mathbb{E}[\hat{U}({\theta}_{a},{\theta}_{b})]$ $:=\mathbb{E}[U({\theta}_{a},{\theta}_{b})]$
$+\sum_{s\in\{a,b\}}n_{s}(u_{+}\alpha_{s}\bm{\Phi}_{s,(i)}^{1}({\theta_{s}})+u_{+}(1-\alpha_{s})\bm{\Phi}_{s,(i)}^{0}({\theta_{s}})+u_{+}\alpha_{s}\bm{\Psi}_{s,(i)}^{1}({\theta_{s}})-u_{-}(1-\alpha_{s})\bm{\Psi}_{s,(i)}^{0}({\theta_{s}})),$ (7)
where the terms $\bm{\Phi}_{s,(i)}^{y}({\theta_{s}})$ (resp. $\bm{\Psi}_{s,(i)}^{y}({\theta_{s}})$) capture the expected value of the change in the firm’s utility relative to the non-strategic setting due to label $y$ agents from group $s$ who opt for improvement (resp. manipulation) when facing classifier $\theta_{s}$ under type (i) best-responses, and are given by
$\bm{\Phi}_{s,(i)}^{y}({\theta_{s}})$ $=\int_{z\in\mathbb{I}^{y}_{s,(i)}}G^{y}_{s}(z)\left(1-\mathbb{T}^{y}_{I,s}({\theta_{s}}-z)\right)\mathrm{d}z,$ (8)
$\bm{\Psi}_{s,(i)}^{y}({\theta_{s}})$ $=\int_{z\in\mathbb{M}^{y}_{s,(i)}}G^{y}_{s}(z)\left(1-\mathbb{T}^{y}_{M,s}({\theta_{s}}-z)\right)\,\mathrm{d}z.$
Note that the first three of these terms increase the firm’s utility over the non-strategic utility: these are all agents who improved and got accepted, and the qualified agents who manipulated their features and got accepted. The last term, with the negative sign, decreases the firm’s utility, as these are unqualified agents who pass the threshold $\theta_{s}$ through manipulation.
Using the above expressions, we can characterize the unfair strategic firm’s optimal policy.
Lemma 5.
The unfair strategic firm’s optimal decision thresholds $\hat{\theta}_{s}^{U}$ for best-response type (i) satisfy $\frac{G_{s}^{1}(\hat{\theta}^{U}_{s})-\bm{\Phi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta}^{U}_{s})-\bm{\Psi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta}^{U}_{s})-\frac{(1-\alpha_{s})}{\alpha_{s}}\bm{\Phi_{s,(i)}^{{}^{\prime}0}}(\hat{\theta}^{U}_{s})}{G_{s}^{0}(\hat{\theta}^{U}_{s})-\bm{\Psi_{s,(i)}^{{}^{\prime}0}}(\hat{\theta}^{U}_{s})}=\frac{u_{-}(1-\alpha_{s})}{u_{+}\alpha_{s}}$, where $\bm{\Phi_{s,(i)}^{{}^{\prime}y}}$ (resp. $\bm{\Psi_{s,(i)}^{{}^{\prime}y}}$) is the first order derivative of $\bm{\Phi_{s,(i)}^{y}}$ (resp. $\bm{\Psi_{s,(i)}^{y}}$) with respect to ${\theta_{s}}$, and with the specific forms under each best-response type given in Table 3.
Table 3: Optimal decision thresholds in Lemma 5.
Type Type 1 Type 2 Type 3
Numerator $C_{I,s}G^{1}_{s}(\mathbf{o}^{1}_{I,s})+(G^{1}_{s}*\tau^{1}_{I,s})(\hat{\theta}^{U}_{s})-(C_{I,s}-C_{M,s})G^{1}_{s}(\mathbf{r}^{1}_{s})+(G^{1}_{s}*\tau^{1}_{M,s})(\hat{\theta}^{U}_{s})+\frac{(1-\alpha_{s})}{\alpha_{s}}\Big{(}C_{I,s}G^{0}_{s}(\mathbf{o}^{0}_{I,s})+(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})-G^{0}_{s}(\mathbf{r}^{0}_{s})\Big{)}$ $C_{M,s}G^{1}_{s}(\mathbf{o}^{1}_{M,s})+(G^{1}_{s}*\tau^{1}_{I,s})(\hat{\theta}^{U}_{s})+(G^{1}_{s}*\tau^{1}_{M,s})(\hat{\theta}^{U}_{s})+G^{1}_{s}(\mathbf{f}^{1}_{s})(C_{I,s}-C_{M,s})-G^{1}_{s}(\mathbf{r}^{1}_{s})(C_{I,s}-C_{M,s})+\frac{(1-\alpha_{s})}{\alpha_{s}}\Big{(}G^{0}_{s}(\mathbf{f}^{0}_{s})(1-\mathbb{T}^{0}_{I,s}(\hat{\theta}_{s}^{U}-\mathbf{f}^{0}_{s}))-G^{0}_{s}(\mathbf{r}^{0}_{s})+(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})\Big{)}$ $C_{M,s}G^{1}_{s}(\mathbf{o}^{1}_{M,s})+(G^{1}_{s}*\tau^{1}_{M,s})(\hat{\theta}^{U}_{s})$
Deno- minator $(1-(C_{I,s}-C_{M,s}))G^{0}_{s}(\mathbf{r}^{0}_{s})+(G^{0}_{s}*\tau^{0}_{M,s})(\hat{\theta}^{U}_{s})$ $C_{M,s}G^{0}_{s}(\mathbf{o}^{0}_{M,s})-(1-\mathbb{T}^{0}_{M,s}(\hat{\theta}_{s}^{U}-\mathbf{f}^{0}_{s}))G^{0}_{s}(\mathbf{f}^{0}_{s})+(1-(C_{I,s}-C_{M,s}))G^{0}_{s}(\mathbf{r}^{0}_{s})+(G^{0}_{s}*\tau^{0}_{M,s})(\hat{\theta}^{U}_{s})$ $C_{M,s}G^{0}_{s}(\mathbf{o}^{0}_{M,s})+(G^{0}_{s}*\tau^{0}_{M,s})(\hat{\theta}^{U}_{s})$
The detailed proof appears in Appendix A.2.2. The main challenge in finding these optimal thresholds is that the regions $\mathbb{{I}}^{y}_{s,(i)}$ and $\mathbb{{M}}^{y}_{s,(i)}$ in which agents opt for the improvement and manipulation actions are functions of the decision variables ${\theta}_{s}$, and these dependencies should be accounted for when evaluating the derivatives $\bm{\Phi_{s,(i)}^{{}^{\prime}y}}$ and $\bm{\Psi_{s,(i)}^{{}^{\prime}y}}$. This is done by applying the Leibniz integral rule, and then leveraging the relation between the threshold and the indifference points in agents’ strategic responses, which appear in the integration limits and their derivatives, to simplify the resulting expressions.
Intuitively, the terms in Table 3 can be interpreted as balancing the marginal gains and losses of a firm at the optimal decision thresholds. To illustrate, let’s take the expression $(1-\alpha_{s})\Big{(}C_{I,s}G^{0}_{s}(\mathbf{o}^{0}_{I,s})+(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})-G^{0}_{s}(\mathbf{r}^{0}_{s})\Big{)}$, appearing in the numerator of Type 1 equilibrium characterization, which reflects the rate of change in the benefits from accepting unqualified agents who opt for improvement, as the decision threshold changes. If the decision threshold $\hat{\theta}^{U}_{s}$ increases by a small $\epsilon$, the firm will lose the agents at $x=\mathbf{o}^{0}_{I,s}$ who successfully made it to the old threshold through improvement (at a rate $(1-\alpha_{s})C_{I,s}G^{0}_{s}(\mathbf{o}^{0}_{I,s})$). The firm will also lose some of those with $x\in(\mathbf{o}^{0}_{I,s},\mathbf{r}^{0}_{I,s})$ who no longer make it to the new threshold (at a rate $(1-\alpha_{s})(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})$), yet will gain from agents with $x=\mathbf{r}^{0}_{s}$ who now opt for improvement instead of manipulation (at a rate $(1-\alpha_{s})G^{0}_{s}(\mathbf{r}^{0}_{s})$). Other expressions can be intuitively interpreted in a similar way.
Lastly, we extend the above two lemmas when the firm also incorporates a fairness constraint in its selection. Lemma 6 characterizes the optimal non-strategic fairness-constrained decision thresholds.
Lemma 6.
The fair non-strategic firm’s optimal decision thresholds $\theta_{s}^{\mathcal{C}}$ satisfy:
$\sum_{s}n_{s}\frac{u_{+}\alpha_{s}G^{1}_{s}(\theta_{s}^{\mathcal{C}})-u_{-}(1-\alpha_{s})G^{0}_{s}(\theta_{s}^{\mathcal{C}})}{\frac{\partial{\mathcal{C}^{f}_{s}(\theta_{s}^{\mathcal{C}})}}{\partial{\theta_{s}^{\mathcal{C}}}}}=0$.
Lemma 7 characterizes the optimal strategic fairness-constrained decision thresholds.
Lemma 7.
The fair strategic firm’s optimal decision thresholds $\hat{\theta}_{s}^{\mathcal{C}}$ satisfy:
$\sum_{s}n_{s}$ $\frac{u_{+}\alpha_{s}(\bm{\Phi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta_{s}^{\mathcal{C}}})+\bm{\Psi_{s,(i)}^{{}^{\prime}1}}(\hat{\theta_{s}^{\mathcal{C}}})-G_{s}^{1}(\hat{\theta_{s}^{\mathcal{C}}}))+(1-\alpha_{s})(u_{+}\bm{\Phi_{s,(i)}^{{}^{\prime}0}}(\hat{\theta_{s}^{\mathcal{C}}})-u_{-}(\bm{\Psi_{s,(i)}^{{}^{\prime}0}(\hat{\theta_{s}^{\mathcal{C}}})}-G_{s}^{0}(\hat{\theta_{s}^{\mathcal{C}}})))}{\frac{\partial{\mathcal{C}^{f}_{s}(\hat{\theta}_{s}^{\mathcal{C}})}}{\partial{\theta_{s}^{\mathcal{C}}}}}=0$.
The proofs follow from applying similar techniques to those in [Zhang et al., 2022], combined with the results of Lemma 5; they are included in Appendix A.2.3 for completeness.
Lemmas 4-7, together with Proposition 2, fully characterize the possible (Stackelberg) equilibria of the game between agents and the firm, depending on the relative cost-efficacy of strategic actions, whether the firm is cognizant of agents’ strategic behavior, and whether the firm imposes any fairness constraints. We next use these characterizations to assess the impacts of anticipating agents’ strategic behavior and imposing fairness constraints on the optimal decision thresholds and the firm’s utility.
6 Effects of Anticipating Strategic Behavior
We first proceed with analyzing the impacts of anticipating agents’ strategic behavior on the optimal policies and the firm’s utility by comparing the non-strategic policy $\theta^{U}_{s}$ (from Lemma 4) with the strategic policy $\hat{\theta}_{s}$ (from Lemma 5) in the absence of any fairness interventions. For concreteness, we do this analysis for cases when all agents opt for the same type of best-response, and specifically, for Types 1 and 3 of agents’ best responses, which can be interpreted as relatively low and high improvement costs, respectively, as illustrated in Section 4.1. Accordingly, we will also refer to Type 1 and Type 3 equilibria as “affordable improvement” and “costly improvement”, respectively. Notably, Type 3/costly improvement equilibria only include manipulation decisions, whereas Type 1/affordable improvement equilibria include both manipulation and improvement decisions at equilibrium. Contrasting these two types of equilibria allows us to compare our findings with prior work, e.g. [Zhang et al., 2022], which have studied the (fairness) implications of anticipating gaming given only manipulation decisions. With our model, we can investigate to what extent anticipating gaming can prevent manipulation (as shown in prior work) but also potentially encourage improvement decisions (as we show here), and how the extent of each differs across qualification states and across groups.
Our main result, Proposition 8, compares the strategic and non-strategic decision thresholds, and their influence on different components of the firm utility. Recall that when a decision threshold $\theta_{s}$ is lowered (resp. increased), more (resp. fewer) agents are accepted without taking any strategic action. The relative ordering of the thresholds therefore allows us to comment on how the strategic responses of agents from different labels and groups changes if a firm can anticipate strategic responses. To establish these results, we make two additional assumptions: one on the feature distributions, and one on the costs and boost functions for different actions across qualification states.
Assumption 3.
Agents’ feature distribution $G^{y}_{s}(x)$ is symmetric and bounded, with lower and upper domain limits $\underline{x}^{y}_{s}$, and $\overline{x}^{y}_{s}$, respectively. Further, $\mu^{0}_{s}<\underline{x}^{1}_{s}<\overline{x}^{0}_{s}<\mu^{1}_{s}$, where $\mu^{y}_{s}$ denotes the mean of $G^{y}_{s}(x)$.
The assumption of symmetric feature distributions is made for tractability of the analysis. The inequality in the assumption poses a natural (and mild) restriction: unqualified agents have, on average, lower scores than qualified agents ($\mu_{s}^{0}<\mu^{1}_{s}$), and there are overlaps in their feature distributions ($\underline{x}^{1}_{s}<\overline{x}^{0}_{s}$) so that distinguishing them based on their feature $x$ is a non-trivial task.
Assumption 4.
The manipulation cost and boost distribution are the same for both labels (i.e., $C^{1}_{M,s}=C^{0}_{M,s}=C_{M,s}$ and $\tau^{1}_{M,s}=\tau^{0}_{M,s}=\tau_{M,s}$). The improvement boost distribution of qualified agents first-order stochastically dominates (FOSD) that of unqualified agents ($i.e.,\mathbb{T}_{I,s}^{0}(b)\geq\mathbb{T}_{I,s}^{1}(b)$).
Assumption 4 states that the cost and effectiveness of the manipulation action is independent of the qualification state of agents. For instance, using a GPT tool to solve basic homework problems has the same cost and effectiveness regardless of a student’s preparation level/qualification state. In contrast, the second part of the assumption implies that improvement action is more effective for qualified agents. Continuing on the same example, when a qualified student with a stronger background spends time preparing well for an exam, their effort can boost their exam grades more than a less prepared student spending the same time.
Given Assumptions 3 and 4, we are now ready to establish relative orderings between $\theta_{s}$ and $\hat{\theta}_{s}$, and their impact on the different components of the firm’s utility. Recall, from the expression for the firm’s utility (5), that $\bm{\Phi}_{s,(i)}^{y}({\theta_{s}})$ and $\bm{\Psi}_{s,(i)}^{y}({\theta_{s}})$ capture the change in the firm’s utility due to label $y$ agents from group $s$ who opt for improvement and manipulation, respectively, in equilibrium Type (i), under a threshold $\theta_{s}$.
Proposition 8.
Assume $\underline{x}^{1}_{s}\geq\mu^{0}_{s}+\underline{b}_{M,s}$, $\overline{x}^{0}_{s}-\underline{x}^{0}_{s}<\mathbf{r}^{y}_{s}-\mathbf{o}^{0}_{I,s}$,$\underline{x}^{1}_{s}-\underline{x}^{0}_{s}>(\mathbb{T}^{y}_{M,s})^{-1}(1-C_{M,s})$, $\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}<\mu^{1}_{s}$, and $\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(1-C_{M,s})<\mu^{1}_{s}+\underline{b}_{M,s}$, and that Assumptions 3 and 4 hold. Then,
  • (i)
    The strategic firm chooses a higher threshold: $\theta^{U}_{s}<\hat{\theta}^{U}_{s}$,
  • (ii)
    The non-strategic firm faces more manipulation by unqualified agents: $\bm{\Psi}^{0}_{s,(i)}(\theta^{U}_{s})>\bm{\Psi}^{0}_{s,(i)}(\hat{\theta}^{U}_{s}),~{}i\in\{1,3\}$,
  • (iii)
    The strategic firm faces more manipulation by qualified agents: $\bm{\Psi}^{1}_{s,(i)}(\theta^{U}_{s})<\bm{\Psi}^{1}_{s,(i)}(\hat{\theta}^{U}_{s}),~{}i\in\{1,3\}$,
  • (iv)
    The strategic firm drives more improvement by both qualified and unqualified agents: $\bm{\Phi}^{y}_{s,(1)}(\theta^{U}_{s})<\bm{\Phi}^{y}_{s,(1)}(\hat{\theta}^{U}_{s}),~{}y\in\{0,1\}$.
The proofs are provided in Appendix A.3. We will next provide intuitive interpretations for the findings of this proposition. We discuss the impacts of relaxing the assumptions of Proposition 8 in Appendix A.4, as well as in the numerical experiments in Section 7.
6.1 Intuitive interpretation and key takeaways of Proposition 8
To highlight the intuition from Proposition 8, we will leverage the illustrations in Figure 2. In this figure, the horizontal, colored bars show the different regions of agents’ strategic best-responses under the costly improvement (Type 3) and affordable improvement (Type 1) equilibria (as shown in Proposition 2), for the non-strategic $\theta^{U}_{s}$ and strategic $\hat{\theta}^{U}_{s}$ policies. The curves above each plot show the pre-strategic feature distribution of agents with qualification state $Y$. The lightly shaded green/red ovals are used to further highlight the agents that opt for improvement/manipulation in each case.
(a) Costly improvement (Type 3) equilibrium
(b) Affordable improvement (Type 1) equilibrium
Figure 2: The effect of the non-strategic and strategic policies on agents’ strategic behavior in different equilibrium types. The curves represent the label $Y$ agents’ feature distributions.
The costly improvement (Type 3) equilibrium.
We start with settings where a Type 3 (manipulation only) equilibrium emerges (e.g., when the improvement cost is too high). A nonstrategic firm, oblivious to the possibility of agents responding strategically, chooses a threshold $\theta^{U}_{s}$ solely with the goal of rejecting unqualified agents; this is seen in the top panel of Figure 2(a). However, as agents are strategic, they will opt for manipulation to increase their chances of surpassing this threshold. Therefore, a non-strategic firm ends up admitting a considerable number of unqualified ($Y=0$) agents (along with a small number of qualified ($Y=1$) agents) who are below the threshold but pass it by opting for manipulation.
A strategic firm accounts for this effect, and chooses a higher $\hat{\theta}^{U}_{s}$; this is asserted in part (i) of Proposition 8. The higher threshold accommodates less manipulation by unqualified ($Y=0$) agents who are now too far from the threshold to benefit from manipulation (as supported by part (ii) of Proposition 8). At the same time, the higher threshold increases the number of qualified ($Y=1$) agents who are no longer accepted by default, and therefore leads to more qualified agents opting for manipulation compared to the non-strategic firm’s case (as seen in part (iii) of Proposition 8); this is also illustrated in the bottom subplot in Figure 2(a). This is however the firm’s desired strategic response by agents, as manipulation by qualified agents, which leads to their acceptance, also benefits the firm and aligns with its objective.
The affordable improvement (Type 1) equilibrium.
We next turn to problem settings where Type 1 (mix of improvement and manipulation) responses emerge, as, e.g., improvement costs are sufficiently low; see Figure 2(b). Here, the strategic firm again chooses a higher threshold than the non-strategic firm. Part of this is for the same reasons outlined earlier: by choosing a higher threshold, the strategic firm takes away manipulation opportunities from unqualified agents (who would otherwise reduce the firm’s utility by getting admitted through manipulation), but increases manipulation by qualified agents (a choice that is non-detrimental to the firm). That said, the more interesting observations, which are new to our model compared to prior work, are the impacts of this increase on the agents’ improvement choices. In the non-strategic firm scenario, a lower threshold leads to only some of the unqualified ($Y=0$) agents opting for improvement, with a significant number opting for manipulation (see top panel of Figure 2(b)). In contrast, the increased threshold set by a strategic firm drives the majority of the unqualified agents to opt for improvement as the only viable option to get admitted (see bottom panel of Figure 2(b)). This is the main driver of the increase in the strategic firm’s utility, and is supported by part (iv) of Proposition 8. We also note that the same observation is true for qualified ($Y=1$) agents.
Together, these observations highlights how anticipating strategic responses will not only reduce gaming of the algorithm (as also shown in prior work), but also lead to increased improvement incentives among agents (a new effect identified by our model).
6.2 Effects of fairness interventions
We now proceed with discussing the effect of imposing a fairness constraint $\mathcal{C}$ when selecting the optimal threshold (as formalized in (3)), on the contrast between the unfair policies (characterized in Lemma 4 and 5) with the corresponding fair policies (characterized in Lemmas 6 and 7). Similar to Section 6, we do so when all agents opt for the same response type, and for Types 1 and 3 of agents’ best responses. Our comparison, outlined in the following Corollary, is a direct consequence of the proof of Proposition 8 under the stated assumptions.
Corollary 9.
For a non-strategic firm, assume that $\theta^{U}_{a}<\theta^{\mathcal{C}}_{a}\leq\arg\max_{\theta_{s}}\bm{\Phi}^{0}_{a,(1)}(\theta_{s})$ and that $\theta^{U}_{b}>\theta^{\mathcal{C}}_{b}\geq\bm{\Phi}^{y}_{b,(1)}(\hat{\theta}^{\mathcal{C}}_{b})$, where $a$ and $b$ denote the advantaged and disadvantaged groups, respectively. Then, imposing a fairness constraint:
  1. 1.
    Decreases (increases) manipulation by unqualified advantaged (disadvantaged) agents ($\bm{\Psi}^{0}_{a,(i)}(\hat{\theta}^{\mathcal{C}}_{a})\leq\bm{\Psi}^{0}_{a,(i)}(\hat{\theta}^{U}_{a})$ and $\bm{\Psi}^{0}_{b,(i)}(\hat{\theta}^{\mathcal{C}}_{b})>\bm{\Psi}^{0}_{b,(i)}(\hat{\theta}^{U}_{b}),~{}i\in\{1,3\}$),
  2. 2.
    Increases (decreases) manipulation by qualified advantaged (disadvantaged) agents (i.e., $\bm{\Psi}^{1}_{a,(i)}(\hat{\theta}^{\mathcal{C}}_{a})\geq\bm{\Psi}^{1}_{a,(i)}(\hat{\theta}^{U}_{a})$ and $\bm{\Psi}^{1}_{b,(i)}(\hat{\theta}^{\mathcal{C}}_{b})\leq\bm{\Psi}^{1}_{b,(i)}(\hat{\theta}^{U}_{b}),~{}i\in\{1,3\}$),
  3. 3.
    Increases (decreases) improvement by both qualified and unqualified advantaged (disadvantaged) agents (i.e., $\bm{\Phi}^{y}_{a,(1)}({\theta}^{\mathcal{C}}_{a})>\bm{\Phi}^{y}_{a,(1)}({\theta}^{U}_{a})$, and $\bm{\Phi}^{y}_{b,(1)}(\hat{\theta}^{\mathcal{C}}_{b})\leq\bm{\Phi}^{y}_{b,(1)}(\hat{\theta}^{U}_{b})$, $y\in\{0,1\}$).
The same statements hold for the strategic firm with the corresponding $\hat{\theta}^{{U}}_{s}$ and $\hat{\theta}^{\mathcal{C}}_{s}$.
These observations are primarily due to the ordering of fair vs. unfair policies, where we have assumed that the firm attains fairness by raising the decision threshold for the advantaged group ($\theta^{\mathcal{C}}_{a}>\theta^{U}_{a}$) and lowering it for the disadvantaged group ($\theta^{\mathcal{C}}_{b}<\theta^{U}_{b}$). This is a natural assumption, supported by our numerical experiments in Section 7, and also commonly observed in existing literature on fair machine learning. In particular, Estornell et al. [2023] discuss that in binary classification, fairness can be improved in two main ways: inclusiveness (flipping disadvantaged group predictions to positive) or selectivity (flipping advantaged group predictions to negative). Our assumption of a decrease in the group $b$ threshold aligns with inclusiveness, while that of an increase in the group $a$ threshold aligns with selectivity, to attain fairness.
In our context, Corollary 9 argues that in its attempt to attain fair outcomes, the firm decreases improvement incentives of the disadvantaged group $b$, while increasing the improvement incentives for the advantaged one; this can be viewed as a negative consequence of fairness interventions on shaping agents’ strategic responses. As we elaborate through numerical experiments in Section 7, this negative by-product is more severe if the firm is non-strategic (i.e., unaware of agents’ strategic responses).
7 Numerical Experiments
In this section, we will use numerical experiments to support our analytical findings and observations from Section 6. We focus on affordable improvement (Type 1) equilibria which contain a mix of manipulation and improvement responses; similar observations can be made for costly improvement (Type 3) equilibria and are provided in Appendix A.5. We will start with a population consisting of a single demographic group to highlight the main observations from Proposition 8 (including under some relaxations of its assumptions).
We will then presents experiments on a population with two distinct demographic groups, placing emphasis on the impacts of fairness interventions, contrasting the induced strategic behavior in each demographic group, and analyzing the induced post-strategic qualification rates across groups at equilibrium.
Single-group experiment setup.
We start with a population of agents from a single group $s$, with truncated Gaussian feature distributions. In particular, we let $G^{0}_{s}\sim\text{TruncNorm}([20,60],40,15^{2})$ and $G^{1}_{s}=\text{TruncNorm}([53,113],83,15^{2})$. We vary the (pre-strategic) qualification rate $\alpha_{s}$ in $(0,1)$. For the characteristics of the strategic actions, we set $C_{M,s}=0.1$, and $C_{I,s}=0.6$. The boost distributions are truncated Gaussian distributions, parameterized as follows: $\tau^{y}_{M,s}(b)\sim\text{TruncNorm}([10,50],30,22^{2}),$, $\tau^{0}_{I,s}(b)\sim\text{TruncNorm}([37,79],58,15^{2})$, and $\tau^{1}_{I,s}(b)\sim\text{TruncNorm}([40,80],60,15^{2}).$ We draw $n=1000$ total agents from the feature distributions, with the number drawn from in each label determined from the selected $\alpha_{s}$. For each $\alpha_{s}$, we run the experiment $50$ times, and report the average across these runs.
Strategic vs. non-strategic (unfair) firms.
In Figure 3, we explore the impacts of anticipating agents’ strategic behavior (in the absence of any fairness interventions). Figure 3(a) shows that the strategic optimal thresholds are greater than the optimal non-strategic threshold at all levels of $\alpha_{s}$; this is consistent with Proposition 8. Figure 3(b) supports this by showing that a strategic firm’s utility is greater than the non-strategic one’s. Lastly, consistent with parts (ii)-(iv) of Proposition 8, we observe that the anticipation of agents’ strategic behavior reduces the impacts of manipulation by the unqualified agents (Fig. 3(c)) while increasing improvement by them (Fig. 3(d)). Simultaneously, this choice leads to both more manipulation (Fig. 3(e)) and (weakly) more improvement (Fig. 3(f)) by the qualified agents, both of which also benefit the firm. Figure 4 further supports these claims, by showing that post-strategic qualification rates $\hat{\alpha}_{s}$ are higher under the strategic firm’s policy compared to the non-strategic firm’s policy. In other words, strategic firms succeed at incentivizing agents to opt for improvement, with the impacts being more pronounced when the group is initially majority unqualified (low $\alpha_{s}$). Finally, we also note that all these observations hold even though the assumption $\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}$ is not satisfied in the current experiment setup, hinting that Proposition 8 may hold beyond its current assumptions.
(a) Optimal thresholds comparison
(b) Optimal utility comparison
(c) $\bm{\Psi}^{0}_{s,(1)}$ (Manipulation) comparison
(d) $\bm{\Phi}^{0}_{s,(1)}$ (Improvement) comparison
(e) $\bm{\Psi}^{1}_{s,(1)}$ (Manipulation) comparison
(f) $\bm{\Phi}^{1}_{s,(1)}$ (Improvement) comparison
Figure 3: Unfair optimal non-strategic vs. strategic policies’ comparison when varying pre-strategic $\alpha_{s}$.
Figure 4: Unfair non-strategic vs. strategic policies’ impact on post-strategic $\hat{\alpha}_{s}$ across pre-strategic $\alpha_{s}$.
Two-group numerical illustration setup.
We next consider a population consisting of two equal-size groups $a$ and $b$, with group $a$ being majority-qualified ($\alpha_{a}=0.7$) and group $b$ being majority-unqualified ($\alpha_{b}=0.2$). The feature distributions for both groups follow truncated Gaussian distributions, with $G^{0}_{a}\sim\text{TruncNorm}([20,110],65,15^{2})$, $G^{0}_{b}\sim\text{TruncNorm}([0,90],45,15^{2})$, $G^{1}_{a}\sim\text{TruncNorm}([98,188],143,15^{2})$, and $G^{1}_{b}\sim\text{TruncNorm}([78,168],123,15^{2})$. We assume both groups have access to the same strategic actions, with $C_{M,s}=0.2$, $C_{I,s}=0.3$, and the boost distributions $\tau^{y}_{M,s}(b)\sim\text{TruncNorm}([20,70],45,22^{2})$, $\tau^{0}_{I,s}(b)\sim\text{TruncNorm}([75,115],95,15^{2})$, and $\tau^{1}_{I,s}(b)\sim\text{TruncNorm}([72,118],97,15^{2})$. Note that with these assumptions, the disparities between groups are assumed to be historical (in qualification rates and feature distributions) but not current (in access to strategic resources).
Strategic vs. non-strategic fair firms.
Figure 5 uses contour plots to illustrate the firms’ utility, and the fairness-constrained utilities under two types of constraints (Demographic Parity and True Positive Rate parity) when varying the group thresholds $\theta_{a}$ and $\theta_{b}$. It also indicates the optimal thresholds for each case (unfair, DP-fair, and TPR-fair). Figure 6 also shows these thresholds, and the manipulation and improvement impacts on the firm’s utility under each choice. It is worth noting that the strategic firm is aware of post-strategic statistics, and uses these when evaluating and imposing the fairness constraints. In contrast, due to unawareness of strategic behavior, the non-strategic firm evaluates its fairness constraints over the pre-strategic constraints, which in turn means that its fairness desiderata are in fact not satisfied ex-post.
We first observe that both strategic and non-strategic firms set higher (resp. lower) thresholds on the majority-qualified (resp. majority-unqualified) group in order to meet a fairness constraint; this is consistent with existing literature and our assumption in Corollary 9. We also note that the strategic firm (whether fair or not) sets higher thresholds on both groups that the non-strategic firm; this is also consistent with Proposition 8.
The more notable and newly observed distinction is when we compare the impact of the type of fairness intervention (DP vs. TPR) on the difference between strategic and non-strategic firms. In particular, there is a more significant change in DP-fair thresholds when the firm is non-strategic. To see why, consider the DP-Fair policy selected by a non-strategic firm (Fig. 5(a)). Such a firm does not account for agents’ strategic responses, simply lowering the decision threshold on group $b$ and increasing it on group $a$, in order to satisfy the fairness constraint. For group $b$, this DP-Fair policy is in fact so low that it accepts many agents (by default) and accepts others with even lower features who opt for manipulation. That is why $\bm{\Phi}^{0}_{b,(1)}(\theta^{\mathcal{C}}_{b})$ (resp. $\bm{\Psi}^{0}_{b,(1)}(\theta^{\mathcal{C}}_{b})$) decreases (resp. increases) compared to those under the unfair non-strategic policy, as illustrated by the solid red lines in Figures 6 the second and the fourth subplots. In contrast, a strategic firm lowers its fair thresholds on group $b$ much less drastically (Fig. 5(b)), as it realizes that parity between selection rates (as required by DP) can be achieved by a combination of adjusting the thresholds and driving agents’ best-responses. In fact, we can observe that fairness desiderata are achieved with very small changes to the decision thresholds when accounting for agents’ strategic responses, therefore leading to far less loss in utility for the firm.
(a) Non-strategic firm’s utility
(b) Strategic firm’s utility
Figure 5: Firm’s utility and optimal thresholds when imposing fairness constraints.
Figure 6: Strategic behavior’s impact comparison: fair vs unfair policies
Impacts on agents’ post-strategic qualification rates.
To further explore the impacts of a strategic policy on agents’ behavior under different fairness interventions (unfair, DP, and TPR), in Figure 7, we contrast the post-strategic $\hat{\alpha}_{s}$ qualification rates for each group. First, as also noted earlier, we note that the strategic firm is more successful at incentivizing improvement, as evidenced by the higher $\hat{\alpha}_{s}$ in both groups $a$ and $b$ compared to the non-strategic firm (in Fig. 7, blue bars are consistently higher than red bars). This increase is even more significant for group $b$, as this group has a lower qualification rate to begin with.
Perhaps more interestingly, we observe that if a firm does not account for strategic responses by agents, fairness interventions can have a negative implication as they decrease improvement by the disadvantaged group $b$ (in Fig. 7, the red bars are lower under the fair policies compared to the unfair policy). This is because fairness interventions typically lower the threshold for group $b$ compared to the unfair policy, which in turn drives the adoption of manipulation over improvement decisions. In comparison, a strategic firm recognizes that fairness goals can be met by a combination of lowering the threshold and incentivizing improvement actions, ultimately leading to higher improvement rates in the disadvantaged group $b$.
Figure 7: Fair vs. unfair, non-strategic vs. strategic policies’ impact on post-strategic $\hat{\alpha}_{s}$.
8 Conclusion and Future Work
We proposed a Stackelberg game model to study strategic classification, where a firm deploys a (fair) classifier and agents strategically respond by adjusting their true qualification states and/or observable features to increase their chances of acceptance. We model both different costs and stochastic efficacy for the agents’ manipulation and improvement actions. We find that anticipating strategic behavior can allow the firm to not only curb manipulation behavior (as also noted in prior work) but can also incentivize agents to opt for improvement decisions. Specifically, we find that a strategic firm chooses its policy to incentivize improvement by unqualified agents (driving them to improve both their qualification states and observable features), while still allowing for manipulation by some qualified agents (who do not have sufficiently high observable features to be selected otherwise). We further show that if a firm chooses to impose a fairness constraint on its decision rule, while remaining unaware of agents’ strategic responses, this choice will typically increase (resp. lower) the thresholds for advantaged (resp. disadvantaged) groups, in turn increasing (resp. decreasing) their incentives for improvement. In terms of its impact on a disadvantaged (majority-unqualified) group, this can be viewed as a negative byproduct of fairness interventions. In contrast, a strategic firm understands that desired fairness goals can be attained through a combination of adjusting the decision thresholds and driving agents’ best-responses, maintaining improvement incentives across all groups.
In conducting our analysis, we held the assumption of similar strategic behavior resources (cost and boost distributions) across groups. A main direction of future work includes exploring the impacts of varying these conditions among groups (e.g., availability of better strategic resources for the advantaged group), and the choice of interventions by the firm (e.g., when the firm makes the improvement action cheaper for a disadvantaged group group by providing free resources that help in improving skills/features), and make it more expensive for other groups. Similarly, this paper focused on studying agents’ strategic behavior when all agents (qualified and unqualified, and regardless of group membership) follow the same type of equilibrium best-responses as characterized in Proposition 2. Varying the resources available to different agents would also lead them to adopt different best-response types It would therefore be interesting to extend our analysis to scenarios where different groups adopt different types of equilibrium responses.
Appendix A Supplementary Material
A.1 Proofs for Section 4
A.1.1 Proof of Proposition 2.
We begin by establishing the possible orderings between the indifference points of Definition 1, under Assumption 2, showing that the cases identified in Proposition 2 cover all possible equilibrium outcomes given the possible orderings of the four constants $\mathbf{o}^{y}_{M,s},\mathbf{o}^{y}_{I,s},\mathbf{f}^{y}_{s},\mathbf{r}^{y}_{s}$.
Lemma 10.
We have $\mathbf{r}^{y}_{s}\geq\{\mathbf{f}^{y}_{s},\mathbf{o}^{y}_{M,s}\}$. Further, if $\mathbf{f}^{y}_{s}$ is unique, either $\mathbf{f}^{y}_{s}\geq\max\{\mathbf{o}^{y}_{M,s},\mathbf{o}^{y}_{I,s}\}$, or $\mathbf{f}^{y}_{s}\leq\min\{\mathbf{o}^{y}_{M,s},\mathbf{o}^{y}_{I,s}\}$. Additionally, if $\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}$, there is at most one $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}$.
Proof.
First, note that $(\mathbb{T}^{y}_{M,s})^{-1}(C_{I}-C_{M,s})\leq(\mathbb{T}^{y}_{M,s})^{-1}(1-C_{M,s})$; this is because the inverse CDF is an increasing function, and $C_{I,s}\leq 1$. Therefore, $\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$. Also, by definition, $\mathbb{T}^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})-\mathbb{T}^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})=\mathbb{T}^{y}_{M,s}(\theta_{s}-\mathbf{r}^{y}_{s})$, and therefore $\mathbf{f}^{y}_{s}\leq\mathbf{r}^{y}_{s}$. Finally, note that if $\mathbf{f}^{y}_{s}\geq\mathbf{o}^{y}_{M,s}$, then $\mathbb{T}^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})=\mathbb{T}^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})-(C_{I,s}-C_{M,s})\leq\mathbb{T}^{y}_{M,s}(\theta_{s}-\mathbf{o}^{y}_{M,s})-(C_{I,s}-C_{M,s})=1-C_{I,s}=\mathbb{T}^{y}_{I,s}(\theta_{s}-\mathbf{o}^{y}_{I,s})$, and therefore $\mathbf{f}^{y}_{s}\geq\mathbf{o}^{y}_{I,s}$. Similarly, we can show that if $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{M,s}$, then $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}$. Lastly, if $\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}$, and given that $u_{s}(0,y,I)<u_{s}(0,y,M)$ due to FOSD, it must be that the utility of $I$ crosses the utility of $M$ at some $x\leq\mathbf{o}^{y}_{I,s}$, and therefore $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}$. ∎
Now, note that if $x\geq\theta_{s}$, the agent is already admitted by the classifier, and finds it optimal to do nothing. As such, only agents with $x<\theta_{s}$ may opt for manipulation or improvement decisions. For these agents, the probability of being admitted if neither action is taken is zero. Together with Eq.(1), this means that the utility of such agents with $x<\theta_{s}$ when choosing $w\in\{M,I\}$ reduces to $u_{s}(x,y,w)=\mathbb{P}(\hat{x}\geq\theta_{s}|X=x,Y=y,W=w,S=s)-C_{w,s}$.
We now proceed by finding the features $\mathbf{o}^{y}_{M,s}$ at which the agent first finds it beneficial to opt for manipulation over doing nothing. Note that the agent’s utility is non-decreasing in $x$. This is because we have assumed adopting either of the two actions $M$ or $I$ weakly increases the agent’s feature, and hence (weakly) increases the probability of being admitted by the classifier. This means that if an agent with feature $\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu$ prefers action $w$ over doing nothing, so will all $x>\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu$.
Recall that $u_{s}(x,y,N)=0$. Therefore, $\mathbf{o}^{y}_{M,s}$ is the first x at which $u_{s}(x,y,M)\geq 0$. This is given by:
$\mathbb{P}(\hat{x}\geq\theta_{s}|X=x,Y=y,W=M,S=s)\geq C_{M,s}\quad\Leftrightarrow\quad x\geq\theta_{s}-(\mathbb{T}^{y}_{M,s})^{-1}(1-C_{M,s})~{}.$
Therefore, the first point at which the agent finds it beneficial to opt for $M$ over $N$ is $\mathbf{o}^{y}_{M,s}=\max\{0,\theta_{s}-(\mathbb{T}^{y}_{M,s})^{-1}(1-C_{M,s})\}$. The first point at which the agent benefits from $I$ over $N$ can be similarly found to be $\mathbf{o}^{y}_{I,s}=\max\{0,\theta_{s}-(\mathbb{T}^{y}_{I,s})^{-1}(1-C_{I,s})\}$. Let $z:=\arg\min_{\{M,I\}}\{\mathbf{o}^{y}_{I,s},\mathbf{o}^{y}_{M,s}\}$. Then, agents with $0\leq x<\min\{\mathbf{o}^{y}_{I,s},\mathbf{o}^{y}_{M,s}\}$ opt for $N$, while those with $\min\{\mathbf{o}^{y}_{I,s},\mathbf{o}^{y}_{M,s}\}\leq x<\max\{\mathbf{o}^{y}_{I,s},\mathbf{o}^{y}_{M,s}\}$ opt for action $z$.
Next, given that the improvement action first-order stochastically dominates the manipulation action, it must be that $\underline{b}^{y}_{M,s}\leq\underline{b}^{y}_{I,s}$ (and also that $\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu^{y}_{M,s}\leq\mkern 1.5mu\overline{\mkern-1.5mub\mkern-1.5mu}\mkern 1.5mu^{y}_{I,s}$). First, we note that once $x\geq\theta_{s}-\underline{b}^{y}_{M,s}$ the agent gets admitted with probability 1 with either $M$ or $I$, and therefore would choose the cheaper action. This means that for $\theta_{s}-\underline{b}^{y}_{M,s}\leq x<\theta_{s}$, the agent chooses action $M$ over $I$. Given the continuity of the utility functions under actions $M$ and $I$, we expect this argument to carry for some $x$ smaller than $\theta_{s}-\underline{b}^{y}_{M,s}$ as well.
Specifically, once $\theta_{s}-\underline{b}^{y}_{I,s}\leq x\leq\theta_{s}-\underline{b}^{y}_{M,s}$, the agent receives sufficient boost to get admitted with probability 1 when choosing action $I$, but is still uncertain when choosing $M$. Formally, improvement has a utility of $1-C_{I,s}$, whereas manipulation has a utility of $1-T^{y}_{M,s}(\theta_{s}-x)-C_{M,s}$. Therefore, if $1-T^{y}_{M,s}(\theta_{s}-x)-C_{M,s}\geq 1-C_{I,s}$ in this region, or equivalently once $x\geq\mathbf{r}^{y}_{s}$, the uncertainty from action $M$ is small enough for the agent to choose action $M$ over $I$. Note also that this argument will continue to hold even if the indifference point $\mathbf{r}^{y}_{s}$ is such that $\mathbf{r}^{y}_{s}\leq\theta_{s}-\underline{b}^{y}_{I,s}$, because this would only increase the uncertainty about $I$, making the utility from choosing $I$ smaller than $1-C_{I,s}$; this means that the utility of action $M$ will still be higher than the utility of $I$ when $x\geq\mathbf{r}^{y}_{s}$, making action $M$ preferable to $I$.
Finally, agents with $\max\{\mathbf{o}^{y}_{I,s},\mathbf{o}^{y}_{M,s}\}\leq x\leq\min\{\mathbf{r}^{y}_{s},\theta_{s}-\underline{b}^{y}_{I,s}\}$ (provided the region is non-empty) would benefit from either manipulation or improvement actions over doing nothing, but face uncertainties about making it to an admit decision when opting for these actions, leading to a cost-efficacy trade off between these choices. Formally, define $\Delta u_{s}(x,y):=u_{s}(x,y,I)-u_{s}(x,y,M)$, the difference between the utility of improvement and manipulation. This difference is:
$\Delta u_{s}(x,y)$ $=\mathbb{P}(\hat{x}\geq\theta_{s}|X=x,Y=y,W=M,S=s)-C_{M,s}$
$\qquad-\left(\mathbb{P}(\hat{x}\geq\theta_{s}|X=x,Y=y,W=I,S=s)-C_{I,s}\right)$
$=\left(\mathbb{T}^{y}_{M,s}(\theta_{s}-x)-\mathbb{T}^{y}_{I,s}(\theta_{s}-x)\right)-\left(C_{I,s}-C_{M,s}\right)~{}.$ (9)
If this difference is positive, the agent will opt for $I$ over $M$. Recall that $\mathbf{f}^{y}_{s}$ is the feature such that $\mathbb{T}^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})-\mathbb{T}^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})=C_{I,s}-C_{M,s}$. This is the point at which the agent is indifferent between the $M$ and $I$ actions. If $M$ is preferred to $I$ before this point, this is the point at which the agent would switch from $M$ to $I$, once $I$ has a non-negative utility (and vice versa for when $I$ is initially preferred to $M$). Using the above characterizations, together with Lemma 10, leads to the identified best responses in each case.
A.1.2 Proof of Lemma 3.
We begin by noting that the new set of qualified ($\hat{Y}=1$) agents consists of previously qualified agents, as well as previously unqualified agents who opted for improvement decisions. Therefore, the new qualification rate is given by $\hat{\alpha}_{s}=\alpha_{s}+(1-\alpha_{s})\int_{\mathbb{I}^{0}_{s}}G^{0}_{s}(x)\mathrm{d}x$.
Next, note that the set of (now) qualified agents with feature ${x}$ consists of the previously qualified agents with the same old feature ${x}$ (who have opted for $w=N$), the previously qualified or unqualified agents with feature $x-b$ who improved to feature $x$ (i.e., opted for $w=I$ and got a boost realization $b$), and previously qualified agents with feature $x-b$ who chose manipulation and reached feature ${x}$ (i.e., opted for $w=M$ and got a boost realization $b$). Thus,
$\hat{G}^{1}_{s}(x)=$ $\tfrac{\alpha_{s}}{\hat{\alpha}_{s}}\big{(}\mathbbm{1}(x\in\mathbb{N}^{1}_{s})G^{1}_{s}(x)+\int_{b}\mathbbm{1}(x-b\in\mathbb{M}^{1}_{s})G^{1}_{s}(x-b)\tau_{M,s}^{1}(b)\mathrm{d}b$
$+\int_{b}\mathbbm{1}(x-b\in\mathbb{I}^{1}_{s})G^{1}_{s}(x-b)\tau_{I,s}^{1}(b)\mathrm{d}b\big{)}+\tfrac{1-\alpha_{s}}{\hat{\alpha}_{s}}\int_{b}\mathbbm{1}(x-b\in\mathbb{I}^{0}_{s})G^{0}_{s}(x-b)\tau_{I,s}^{0}(b)\mathrm{d}b.$ (10)
We can re-write the expressions for the integrals using a change of variable $z:=x-b$ as follows:
$\int_{0}^{\infty}\mathbbm{1}(x-b\in\mathbb{M}^{1}_{s})G^{1}_{s}(x-b)\tau_{M,s}^{1}(b)\mathrm{d}b=\int_{x}^{-\infty}\mathbbm{1}(z\in\mathbb{M}^{1}_{s})G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}(x-z)$
$=$ $\int_{-\infty}^{x}\mathbbm{1}(z\in\mathbb{M}^{1}_{s})G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}z\int_{z\in\mathbb{M}^{1}_{s}}G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}z~{}.$
Substituting the above in Eq.(10) leads to Eq.(6).
Using similar arguments, the post-strategic feature distribution of (now) unqualified agents from group $s$ consists of unqualified agents with the same feature who have opted to do nothing, and previously unqualified agents who have opted for manipulation and have reached feature $x$. Thus:
$\hat{G}^{0}_{s}(x)=$ $\tfrac{1-\alpha_{s}}{1-\hat{\alpha}_{s}}\big{(}\mathbbm{1}(x\in\mathbb{N}^{0}_{s})G^{0}_{s}(x)+\int_{b}\mathbbm{1}(x-b\in\mathbb{M}^{0}_{s})G^{0}_{s}(x-b)\tau_{M,s}^{0}(b)\mathrm{d}b\big{)}$ (11)
Re-writing the integral similar to before leads to Eq.(5).
A.2 Proofs for Section 5
A.2.1 Proof of Lemma 4.
The proof is similar to those in Zhang et al. [2022], Liao and Naghizadeh [2022]. The unfair non-strategic firm’s utility Eq.(2), and its first derivative, are given by
$U(\theta^{U}_{a},\theta^{U}_{b})$ $=\sum_{s}~{}u_{+}\alpha_{s}(1-\mathbb{G}^{1}_{s}(\theta^{U}_{s}))-u_{-}(1-\alpha_{s})(1-\mathbb{G}^{0}_{s}(\theta^{U}_{s})).$
$\Rightarrow\quad\frac{\partial{U(\theta^{U}_{s}})}{\partial{\theta^{U}_{s}}}$ $=-u_{+}\alpha_{s}G^{1}_{s}(\theta^{U}_{s})+u_{-}(1-\alpha_{s})G^{0}_{s}(\theta^{U}_{s}).$
where $\mathbb{G}^{y}_{s}$ denotes the CDF of the feature distribution of agents with true label $y$ from group $s$. Taking the second derivative, and considering Assumption 1, we can see that the firm’s utility is convex in $\theta_{s}$. Therefore, by setting the first derivative to zero, the optimal thresholds satisfy $\frac{G^{1}_{s}(\theta^{U}_{s})}{G^{0}_{s}(\theta^{U}_{s})}=\frac{u_{-}(1-\alpha_{s})}{u_{+}\alpha_{s}}$.
A.2.2 Proof of Lemma 5.
From the unfair strategic firm’s utility Eq.(5) can be written as
$\hat{U}(\hat{\theta}^{U}_{s})$ $=\sum_{s}~{}~{}u_{+}\alpha_{s}(1-\mathbb{G}^{1}_{s}(\hat{\theta}^{U}_{s}))-u_{-}(1-\alpha_{s})(1-\mathbb{G}^{0}_{s}(\hat{\theta}^{U}_{s}))+u_{+}\alpha_{s}\bm{\Phi}^{1}_{s,(i)}(\hat{\theta}^{U}_{s})+u_{+}\alpha_{s}\bm{\Psi}^{1}_{s,(i)}(\hat{\theta}^{U}_{s})$
$+u_{+}(1-\alpha_{s})\bm{\Phi}^{0}_{s,(i)}(\hat{\theta}^{U}_{s})-u_{-}(1-\alpha_{s})\bm{\Psi}^{0}_{s,(i)}(\hat{\theta}^{U}_{s}).$
$\frac{\partial{\hat{U}(\hat{\theta}^{U}_{s}})}{\partial{\hat{\theta}^{U}_{s}}}$ $=-u_{+}\alpha_{s}G^{1}_{s}(\hat{\theta}^{U}_{s})+u_{-}(1-\alpha_{s})G^{0}_{s}(\hat{\theta}^{U}_{s})+u_{+}\alpha_{s}\bm{\Phi}_{s,(i)}^{{}^{\prime}1}+u_{+}\alpha_{s}\bm{\Psi}_{s,(i)}^{{}^{\prime}1}$
$+u_{+}(1-\alpha_{s})\bm{\Phi}_{s,(i)}^{{}^{\prime}0}-u_{-}(1-\alpha_{s})\bm{\Psi}_{s,(i)}^{{}^{\prime}0}.$
where $\bm{\Phi}_{s,(i)}^{{}^{\prime}y}$ and $\bm{\Psi}_{s,(i)}^{{}^{\prime}y}$ are the first derivative of the improvement impact and manipulation impact for agents Y=y at $\hat{\theta}^{U}_{s}$, respectively. Therefore, by setting the above equation of the first derivative of the unfair strategic firm’s utility to zero we have $\frac{\bm{\Phi_{s,(i)}^{{}^{\prime}1}}+\bm{\Psi_{s,(i)}^{{}^{\prime}1}}-G_{s}^{1}(\hat{\theta}^{U}_{s})+\frac{(1-\alpha_{s})}{\alpha_{s}}\bm{\Phi_{s,(i)}^{{}^{\prime}0}}}{\bm{\Psi_{s,(i)}^{{}^{\prime}0}}-G_{s}^{0}(\hat{\theta}^{U}_{s})}=\frac{u_{-}(1-\alpha_{s})}{u_{+}\alpha_{s}}.$
It remains to find the specific forms of the first derivatives $\bm{\Phi}_{s,(i)}^{{}^{\prime}y}$ and $\bm{\Psi}_{s,(i)}^{{}^{\prime}y}$ for each equilibrium type (i).
We show this for agents’ best-response Type 1, finding the expressions in the first row in Table 3. A similar procedure can be used to find the unfair strategic decision threshold for agents’ best-response Type 2 and 3.
From Proposition 2, we know that the improvement set $\mathbb{I}^{y}_{s,(1)}$ is bounded by $\mathbf{o}^{y}_{I,s}$ and $\mathbf{r}^{y}_{s}$ for Type 1 best-response, while the manipulation set $\mathbb{M}^{y}_{s,(i)}$ is bounded By $\mathbf{r}^{y}_{s}$ and $\theta_{s}$. Therefore, we have
$\bm{\Phi}_{s,(1)}^{y}({\theta_{s}})$ $=\int_{\mathbf{o}^{y}_{I,s}}^{\mathbf{r}^{y}_{s}}G^{y}_{s}(z)\left(1-\mathbb{T}^{y}_{I,s}({\theta_{s}}-z)\right)\mathrm{d}z,~{}\text{and }\bm{\Psi}_{s,(1)}^{y}({\theta_{s}})$ $=\int_{\mathbf{r}^{y}_{s}}^{\theta_{s}}G^{y}_{s}(z)\left(1-\mathbb{T}^{y}_{M,s}({\theta_{s}}-z)\right)\,\mathrm{d}z.$
We next find the first derivative of these impacts with respect to $\theta_{s}$. Note that all the limits of the integration are functions of $\theta_{s}$; hence, we followed the Leibniz integral rule to find these derivatives. We also note that the derivatives of $\mathbf{o}^{y}_{I,s}$ and $\mathbf{r}^{y}_{s}$ with respect to $\theta$ are 1. For the improvement impact, we have:
$\bm{\Phi}_{s,(1)}^{{}^{\prime}y}({\theta_{s}})={G}^{y}_{s}(\mathbf{r}^{y}_{s})\left(1-\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{r}^{y}_{s})\right)-{G}^{y}_{s}(\mathbf{o}^{y}_{I,s})\left(1-\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{o}^{y}_{I,s})\right)-\int_{\mathbf{o}^{y}_{I,s}}^{\mathbf{r}^{y}_{s}}G^{y}_{s}(z)\tau^{y}_{I}({\theta_{s}}-z)\mathrm{d}z).$
From the definition of the indifference points in Definition 1, we have $\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{o}^{y}_{I,s})=\mathbb{T}^{y}_{I,s}((\mathbb{T}^{y}_{I,s})^{-1}(1-C_{I,s}))=1-C_{I,s}$. We also know, from Proposition 2, that the conditions for this equilibrium to occur are that $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$, and therefore, $11-\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{r}^{y}_{s})=$ 1 (i.e., admission under $I$ is certain at $\mathbf{r}^{y}_{s}$). Therefore:
$\bm{\Phi}_{s,(1)}^{{}^{\prime}y}({\theta_{s}})={G}^{y}_{s}(\mathbf{r}^{y}_{s})-{G}^{y}_{s}(\mathbf{o}^{y}_{I,s})C_{I,s}-(G^{y}_{s}*\tau^{y}_{I,s})({\theta_{s}}).$ (12)
Similarly, for the manipulation impact we have,
$\bm{\Psi}_{s,(1)}^{{}^{\prime}y}({\theta_{s}})={G}^{y}_{s}(\theta_{s})\left(1-\mathbb{T}^{y}_{M,s}({\theta_{s}}-{\theta_{s}})\right)-{G}^{y}_{s}(\mathbf{r}^{y}_{s})\left(1-\mathbb{T}^{y}_{M,s}({\theta_{s}}-\mathbf{r}^{y}_{s})\right)-\int_{\mathbf{r}^{y}_{s}}^{\theta_{s}}G^{y}_{s}(z)\tau^{y}_{M,s}({\theta_{s}}-z)\mathrm{d}z).$
From the definition of the indifference points in Definition 1, we have $\mathbb{T}^{y}_{M,s}({\theta_{s}}-\mathbf{r}^{y}_{s})=\mathbb{T}^{y}_{M,s}((\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s}))=C_{I,s}-C_{M,s}$. Therefore,
$\bm{\Psi}_{s,(1)}^{{}^{\prime}y}({\theta_{s}})={G}^{y}_{s}(\theta_{s})-{G}^{y}_{s}(\mathbf{r}^{y}_{s})(1-(C_{I,s}-C_{M,s}))-(G^{y}_{s}*\tau^{y}_{M,s})({\theta_{s}}).$ (13)
By substituting these in the fraction determining the optimal unfair strategic threshold in the statement of Lemma 5, we can find the expressions given in Table 3. Note that the derivative of all indifference points defined in Proposition 2 with respect to the decision threshold $\theta_{s}$ is 1, specifically, for the flip feature $(\mathbf{f}^{y}_{s})$ which can be expressed as $\mathbf{f}^{y}_{s}=\theta_{s}-(\mathbb{T}^{y}_{M,s})^{-1}\big{(}\mathbb{T}^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})+C_{I,s}-C_{M,s}\big{)}.$ Let us define: $g(\theta_{s},\mathbf{f}^{y}_{s})=\mathbb{T}^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})+C_{I,s}-C_{M,s},$ and $h(g(\theta_{s},\mathbf{f}^{y}_{s}))=(\mathbb{T}^{y}_{M,s})^{-1}(g(\theta_{s},\mathbf{f}^{y}_{s})).$ Then, the derivative of $\mathbf{f}^{y}_{s}$ with respect to $\theta_{s}$ can be written as:
$\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}=1-\frac{\partial h(g(\theta_{s},\mathbf{f}^{y}_{s}))}{\partial\theta_{s}}.\text{ Next, the derivative of $h(g(\theta_{s},\mathbf{f}^{y}_{s}))$ with respect to $\theta_{s}$ is given by:}$:
$\frac{\partial h(g(\theta_{s},\mathbf{f}^{y}_{s}))}{\partial\theta_{s}}$ $=\frac{1}{\tau^{y}_{M,s}\big{(}(\mathbb{T}^{y}_{M,s})^{-1}(g(\theta_{s},\mathbf{f}^{y}_{s}))\big{)}}\cdot\frac{\partial g(\theta_{s},\mathbf{f}^{y}_{s})}{\partial\theta_{s}}.~{}=\frac{\tau^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})}{\tau^{y}_{M,s}((\mathbb{T}^{y}_{M,s})^{-1}(g(\theta_{s},\mathbf{f}^{y}_{s})))}\cdot\big{(}1-\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}\big{)}.$
Simplifying further, and noting that $\tau^{y}_{M,s}((\mathbb{T}^{y}_{M,s})^{-1}(g(\theta_{s},\mathbf{f}^{y}_{s})))=\tau^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})$, we have:
$\frac{\partial h(g(\theta_{s},\mathbf{f}^{y}_{s}))}{\partial\theta_{s}}=\frac{\tau^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})}{\tau^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})}\cdot\big{(}1-\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}\big{)}.\text{ Substituting this back into the expression for $\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}$},\text{ we get:}$:
$\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}=1-\frac{\tau^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})}{\tau^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})}\cdot\big{(}1-\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}\big{)}.\text{ Rearranging terms yields:}\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}=1.$
A.2.3 Proof of Lemmas 6-7.
Similar to Zhang et al. [2022], from the non-strategic firm’s utility with the fairness constraint in Eq.3 the optimal threshold satisfies $\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})=\mathcal{C}^{f}_{b}(\theta^{\mathcal{C}}_{b})$, thus we can write one of the groups’ fair decision thresholds as $\theta^{\mathcal{C}}_{b}=(\mathcal{C}^{f}_{b})^{-1}(\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a}))$. The following holds:
$\frac{\partial{{\theta}^{\mathcal{C}}_{b}}}{\partial{{\theta}^{\mathcal{C}}_{a}}}=\frac{\partial{(\mathcal{C}^{f}_{b})^{-1}(\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a}))}}{\partial{{\theta}^{\mathcal{C}}_{a}}}=\frac{\partial{(\mathcal{C}^{f}_{b})^{-1}(\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a}))}}{\partial{{\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})}}}\frac{\partial{\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})}}{\partial{{\theta}^{\mathcal{C}}_{a}}}$ $=\frac{1}{(\mathcal{C}^{{}^{\prime}f}_{b})\left((\mathcal{C}^{f}_{b})^{-1}(\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a}))\right)}\frac{\partial{\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})}}{\partial{{\theta}^{\mathcal{C}}_{a}}}$
$=\frac{1}{(\mathcal{C}^{{}^{\prime}f}_{b})\left(\theta^{\mathcal{C}}_{b}\right)}\frac{\partial{\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})}}{\partial{{\theta}^{\mathcal{C}}_{a}}}=\frac{\frac{\partial{\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})}}{\partial{{\theta}^{\mathcal{C}}_{a}}}}{\frac{\partial{\mathcal{C}^{f}_{b}(\theta^{\mathcal{C}}_{b})}}{\partial{{\theta}^{\mathcal{C}}_{b}}}}.$
then the firms’ utility can be written in terms of $\theta^{\mathcal{C}}_{s}$ as,
$U(\theta^{\mathcal{C}}_{a},\theta^{\mathcal{C}}_{b})$ $=n_{a}\left(u_{+}\alpha_{a}(1-\mathbb{G}^{1}_{a}(\theta^{\mathcal{C}}_{a}))-u_{-}(1-\alpha_{a})(1-\mathbb{G}^{0}_{a}(\theta^{\mathcal{C}}_{a}))\right)$
$+n_{b}\left(u_{+}\alpha_{b}(1-\mathbb{G}^{1}_{b}((\mathcal{C}^{f}_{b})^{-1}(\mathcal{C}^{f}_{a}(\theta_{a}))))-u_{-}(1-\alpha_{b})(1-\mathbb{G}^{0}_{b}((\mathcal{C}^{f}_{b})^{-1}(\mathcal{C}^{f}_{a}(\theta_{a}))))\right).$
$\frac{\partial{U(\theta^{\mathcal{C}}_{a},\theta^{\mathcal{C}}_{b})}}{\partial{\theta^{\mathcal{C}}_{a}}}$ $=n_{a}\left(-u_{+}\alpha_{a}G^{1}_{a}(\theta^{\mathcal{C}}_{a})+u_{-}(1-\alpha_{a})G^{0}_{s}(\theta^{\mathcal{C}}_{a})\right)$
$+n_{b}\left(-u_{+}\alpha_{b}G^{1}_{b}(\theta^{\mathcal{C}}_{b}))\frac{\frac{\partial{\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})}}{\partial{{\theta}^{\mathcal{C}}_{a}}}}{\frac{\partial{\mathcal{C}^{f}_{b}(\theta^{\mathcal{C}}_{b})}}{\partial{{\theta}^{\mathcal{C}}_{b}}}}+u_{-}(1-\alpha_{b})G^{0}_{b}(\theta^{\mathcal{C}}_{b})\frac{\frac{\partial{\mathcal{C}^{f}_{a}(\theta^{\mathcal{C}}_{a})}}{\partial{{\theta}^{\mathcal{C}}_{a}}}}{\frac{\partial{\mathcal{C}^{f}_{b}(\theta^{\mathcal{C}}_{b})}}{\partial{{\theta}^{\mathcal{C}}_{b}}}}\right).$
By setting the first derivative to zero and re-arrange terms we have $\theta_{s}^{\mathcal{C}}$ satisfy: $\sum_{s}n_{s}\frac{u_{+}\alpha_{s}G^{1}_{s}(\theta_{s}^{\mathcal{C}})-u_{-}(1-\alpha_{s})G^{0}_{s}(\theta_{s}^{\mathcal{C}})}{\frac{\partial{\mathcal{C}^{f}_{s}(\theta_{s}^{\mathcal{C}})}}{\partial{\theta_{s}^{\mathcal{C}}}}}=0$.
A similar approach can prove Lemma 7 using the strategic utility from Eq. 5. We omit it due to space constraints.
A.3 Proof of Proposition 8
For this part of the proof, we analyze the selected optimal unfair thresholds ($\theta^{U}_{s},\hat{\theta}^{U}_{s}$) over different groups’ qualification status ($\alpha_{s}$) and agents’ best response types (Type 1 and 3). Let $\alpha_{s}=0$ such that all agents are unqualified, the unfair non-strategic firm’s utility in equation 2 can be reduced to $U(\theta_{s})=-u_{-}(1-\mathbb{G}^{0}_{s}(\theta_{s}))$, hence to achieve the maximum utility the optimal threshold is $\theta_{s}=\overline{x}^{0}_{s}$ which means rejecting all the unqualified agents. Moreover, Let $\alpha_{s}=1$ such that all agents are qualified, the unfair non-strategic firm’s utility in equation 2 can be reduced to $U(\theta_{s})=u_{+}(1-\mathbb{G}^{1}_{s}(\theta_{s}))$, hence the optimal threshold is $\theta_{s}=\underline{x}^{1}_{s}$ which means accepting majority all qualified agents. Note as the group qualification rate ($\alpha_{s}$) decreases, the unfair non-strategic optimal threshold increases to reject the more unqualified agents. To summarize, the unfair non-strategic optimal policy satisfies the following: $[\underline{x}^{1}_{s},\overline{x}^{0}_{s}]$. Under Assumption 4, $\mathbf{o}^{y}_{M,s}$ and $\mathbf{r}^{y}_{s}$ are equal across agents, so we denote them as $\mathbf{o}_{M,s}$ and $\mathbf{r}_{s}$, omitting the $y$ label in this proof. To establish a unique optimal strategic utility under Type 3 agent responses, we analyze the derivative from Table 3:
$\frac{\partial{\hat{U}(\hat{\theta}^{U}_{s})}}{\partial{\hat{\theta}^{U}_{s}}}=u_{+}\alpha_{s}\left(-G^{1}_{s}(\mathbf{o}_{M,s})C_{M,s}-(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right)+u_{-}(1-\alpha_{s})\left(G^{0}_{s}(\mathbf{o}_{M,s})C_{M,s}+(G^{0}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right).$
For $\hat{\theta}^{U}_{s}\leq\mu^{0}_{s}$, the derivative is positive, led by $(G^{0}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})$. Per Proposition 8 ($\underline{x}^{1}_{s}-\underline{x}^{0}_{s}>(\mathbb{T}^{y}_{M,s})^{-1}(1-C_{M,s})$), $\mathbf{o}_{M,s}>\underline{x}^{0}_{s}$ before $\hat{\theta}^{U}_{s}>\underline{x}^{1}_{s}$, so $G^{0}_{s}(\mathbf{o}_{M,s})$ rises, boosting the derivative as $\hat{\theta}^{U}_{s}$ nears $\underline{x}^{1}_{s}$. Past $\underline{x}^{1}_{s}$, it drops as $(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})$ grows, possibly turning negative or not. The fall persists as $\mathbf{o}_{M,s}>\mu^{0}_{s}$ approaches $\underline{x}^{1}_{s}$. At $\mathbf{o}_{M,s}>\overline{x}^{0}_{s}$, only negative terms $\left(-G^{1}_{s}(\mathbf{o}_{M,s})C_{M,s}-(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right)$ remain, forcing the derivative negative. As $\hat{\theta}^{U}_{s}\geq\overline{x}^{1}_{s}$, the $(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})$ shrinks, lifting the derivative until $\mathbf{o}_{M,s}>\overline{x}^{1}_{s}$, where it reaches zero, proving a unique optimum.
Next, we bound the convolution terms by treating the convolution as a weighted average of $G^{y}_{s}(x)$ over the interval $\mathbb{W}^{y}_{s}$, with weights $\tau^{y}_{w,s}(\hat{\theta}^{U}_{s}-x)$. The upper bound is,
$(G^{y}_{s}*\tau_{w,s})(\hat{\theta}^{U}_{s})\leq G^{y}_{s}(z_{max})\int_{u\in\mathbb{W}^{y}_{s}}\tau^{y}_{w,s}(\hat{\theta}^{U}_{s}-u)\mathrm{d}u,$
where $z_{max}=\arg\max_{z\in\mathbb{W}^{y}_{s}}G^{y}_{s}(z)$. Furthermore, the lower bound is,
$(G^{y}_{s}*\tau_{w,s})(\hat{\theta}^{U}_{s})\geq G^{y}_{s}(z_{min})\int_{u\in\mathbb{W}^{y}_{s}}\tau^{y}_{w,s}(\hat{\theta}^{U}_{s}-u)\mathrm{d}u,$
where $z_{min}=\arg\min_{z\in\mathbb{W}^{y}_{s}}G^{y}_{s}(z)$. For instance, under Type 1 equilibrium $\mathbb{I}^{0}_{s}=[\mathbf{o}^{0}_{I,s},\mathbf{r}_{s}]$, if $\mathbf{o}^{0}_{I,s}<\mathbf{r}_{s}<\mu^{0}_{s}$ then the upper bound $(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})\leq G^{0}_{s}(\mathbf{r}_{s})(1-C_{I,s})$. Moreover, $\mathbb{M}^{1}_{s}=[\mathbf{r}_{s},\hat{\theta}^{U}_{s}]$, if $\mathbf{r}_{s}<\mu^{1}_{s}$ the lower bound $(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\geq G^{1}_{s}(\mathbf{r}_{s})(C_{I,s}-C_{M,s})$.
For Type 1 agent best responses, the firms’ utility first derivative from Table 3 is:
$\frac{\partial{\hat{U}(\hat{\theta}^{U}_{s})}}{\partial{\hat{\theta}^{U}_{s}}}=$ $u_{+}\alpha_{s}\left(G^{1}_{s}(\mathbf{r}_{s})(C_{I,s}-C_{M,s})-G^{1}_{s}(\mathbf{o}^{1}_{I,s})C_{I,s}-(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})-(G^{1}_{s}*\tau^{1}_{I,s})(\hat{\theta}^{U}_{s})\right)$
$+u_{+}(1-\alpha_{s})\left(G^{0}_{s}(\mathbf{r}_{s})-G^{0}_{s}(\mathbf{o}^{0}_{I,s})C_{I,s}-(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})\right)$
$+u_{-}(1-\alpha_{s})\left(G^{0}_{s}(\mathbf{r}_{s})(1-(C_{I,s}-C_{M,s}))+(G^{0}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right).$
For $\hat{\theta}^{U}_{s}<\underline{x}^{0}_{s}+(\mathbb{T}_{M,s})^{-1}(C{I,s}-C_{M,s})$, the derivative is positive, driven by $(G^{0}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})$. It rises as $\hat{\theta}^{U}_{s}$ nears $\mu^{0}_{s}+(\mathbb{T}_{M,s})^{-1}(C_{I,s}-C_{M,s})$ due to increasing $G^{0}_{s}(\mathbf{r}_{s})$, and per convolution bounds, $(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})\leq G^{0}_{s}(\mathbf{r}_{s})(1-C_{I,s})<G^{0}_{s}(\mathbf{r}_{s})$. Near $\underline{x}^{1}_{s}$, the derivative drops as $\mathbf{r}_{s}>\mu^{0}_{s}$ (positive terms shrink, while negative $(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})$ grows with $\mathbb{I}^{0}_{s}$), possibly staying positive or turning negative. Beyond $\underline{x}^{1}_{s}$, it keeps decreasing as $(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})>0$ grows. When $\underline{x}^{1}_{s}\leq\mathbf{r}_{s}\leq\mu^{1}_{s}$, bounds shows $(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\geq G^{1}_{s}(\mathbf{r}_{s})(C_{I,s}-C_{M,s})$ and $(G^{1}_{s}*\tau^{1}_{I,s})(\hat{\theta}^{U}_{s})>0$, ensuring a decline. Given Proposition 8 ($\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}$), $\mu^{1}_{s}>\mathbf{r}_{s}>\overline{x}^{0}_{s}$ before $\mathbf{o}^{0}_{I,s}>\mu^{0}_{s}$, so $G^{0}_{s}(\mathbf{r}_{s})$ and $(G^{0}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})$ hits zero, turning the derivative negative. As $\mathbf{o}^{0}_{I,s}$ approaching $\mu^{0}_{s}$, only negative terms persist, driving further decrease. When $\mathbf{o}^{0}_{I,s}>\mu^{0}_{s}$, $\left(-G^{0}_{s}(\mathbf{o}^{0}_{I,s})C_{I,s}-(G^{0}_{s}*\tau^{0}_{I,s})(\hat{\theta}^{U}_{s})\right)$ shrinks, increasing the derivative, yet it stays negative since $(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\geq G^{1}_{s}(\mathbf{r}_{s})(C_{I,s}-C_{M,s})$. As $\hat{\theta}^{U}_{s}$ rises, the derivative increases, still negative, with $\mathbf{o}^{1}_{I,s}<\mu^{1}_{s}<\mathbf{r}_{s}$, and bound $(G^{1}_{s}*\tau^{1}_{I,s})(\hat{\theta}^{U}_{s})\geq G^{1}_{s}(\mu^{1}_{s})(1-C_{I,s})\geq G^{1}_{s}(\mathbf{r}_{s})(1-C_{I,s})$. Then, as $G^{0}_{s}(\mathbf{o}^{0}_{I,s})=0$ while $\mathbf{o}^{1}_{I,s}<\mu^{1}_{s}$, causing the derivative to decrease again. As $G^{1}_{s}(\mathbf{r}_{s})=0$ and $\mathbf{o}^{1}_{I,s}$ nears $\mu^{1}_{s}$, it continues decreasing until $\mathbf{o}^{1}_{I,s}>\mu^{1}_{s}$, then rises until $G^{1}_{s}(\mathbf{o}^{1}_{I,s})=0$, hitting zero, confirming a unique optimum.
Under Type 1 agent best responses, when $\alpha_{s}=0$, the unfair strategic firm’s utility (Eq. 5) simplifies to $\hat{U}(\hat{\theta}^{U}_{s})=u_{+}\Phi^{0}_{s,(1)}(\hat{\theta}^{U}_{s})-u_{-}(1-\mathbb{G}^{0}_{s}(\hat{\theta}^{U}_{s}))-u_{-}\Psi^{0}_{s,(1)}(\hat{\theta}^{U}_{s})$. At $\hat{\theta}^{U}_{s}=\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s})$, it is $\hat{U}(\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s}))=u_{+}\Phi^{0}_{s,(1)}(\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s}))$. Notably, utility decreases for $\hat{\theta}^{U}_{s}>\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s})$, where $|\mathbb{I}^{0}_{s,(1)}|$ declines as $\mathbf{r}^{y}_{s}>\overline{x}^{0}_{s}$. Thus, the optimal unfair strategic threshold $\hat{\theta}^{U}_{s}\leq\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s})$
Next we show that the impacts of the strategic behavior are single peaked. For a fixed value of $\theta_{s}$, let $h^{y}_{M,s}(z;\theta_{s}):=G^{y}_{s}(z)\left(1-\mathbb{T}^{y}_{M,s}(\theta_{s}-z)\right).$ The function $h^{y}_{M,s}(z;\theta)$ is single-peaked at $z^{*}$ because $G^{y}_{s}(z)$ is symmetric, and $\left(1-\mathbb{T}^{y}_{M,s}(\theta-z)\right)\in[0,1]$. Moreover,
$\left(1-\mathbb{T}^{y}_{M,s}(\theta-z^{\prime})\right)=\begin{cases}1,&\text{if }z^{\prime}\geq\theta-\underline{b}^{y}_{M,s},\\ 0,&\text{if }z^{\prime}\leq\theta-\overline{b}^{y}_{M,s},\\ (0,1),&\text{otherwise}.\end{cases}$
We can re-write $\bm{\Psi}^{y}_{s,(3)}(\theta_{s})=\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta)\mathrm{d}z$, and because the integration limits depends on $\theta_{s}$, the following holds: $\text{If }\mathbf{o}_{M,s}<\theta_{s}\leq z^{*},\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta_{s})\mathrm{d}z\text{ increases as $\theta_{s}$ approaches the $z^{*}$},$ and $\text{If }z^{*}\leq\mathbf{o}_{M,s}<\theta_{s},\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta_{s})\mathrm{d}z\text{ decreases as $\mathbf{o}_{M,s}$ moves away from the $z^{*}$}.$ Therefore, $\bm{\Psi}^{y}_{s,(3)}(\theta_{s})$ is also single-peaked. Similarly, $\bm{\Psi}^{y}_{s,(i)}(\theta_{s})\text{ for }i\in\{1,2,3\}$, and $\bm{\Phi}^{y}_{s,(i)}(\theta_{s})\text{ for }i\in\{1,2\}$ are also single peaked.
We define conditions for the peak values of $\bm{\Phi}^{y}_{s,(i)}(\theta_{s})$ and $\bm{\Psi}^{y}_{s,(i)}(\theta_{s})$, $y\in{0,1}$. For $\arg\max_{\theta_{s}}\bm{\Psi}^{y}_{s,(i)}(\theta_{s})$, $i\in{1,3}$, $\theta_{s}\leq\mu^{y}_{s}+\underline{b}_{M,s}$ and $\mathbf{r}^{y}_{s}<\mu^{y}_{s}$ ensure agents at $\mu^{y}_{s}$ manipulate and are accepted with minimal boost $\underline{b}_{M,s}$. For $\arg\max_{\theta_{s}}\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$, $\theta_{s}\leq\mu^{y}_{s}+\underline{b}^{y}_{I,s}$ and $\mathbf{r}^{y}_{s}>\mu^{y}_{s}$ compel agents at $\mu^{y}_{s}$ to improve and gain acceptance with minimal boost. Given $\overline{x}^{0}_{s}-\underline{x}^{0}_{s}<\mathbf{r}^{y}_{s}-\mathbf{o}^{0}_{I,s}$, if $\theta_{s}>\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})$ ($\mathbf{r}^{y}_{s}>\overline{x}^{0}_{s}$), $|\mathbb{I}^{0}_{s,(1)}|$ remains unchanged, so $\arg\max_{\theta_{s}}\bm{\Phi}^{0}_{s,(1)}(\theta_{s})\in[\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s}),\mu^{0}_{s}+\underline{b}^{0}_{I,s}]$.
Given the analysis of optimal thresholds, we first order the policies as in Proposition 8 and then rank their impacts as per parts (i), (ii), and (iii) of the same proposition. The maximum value of the optimal unfair strategic threshold, $\hat{\theta}^{U}_{s}$, under Type 1 agent responses is $\hat{\theta}^{U}_{s}\leq\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}<\mu^{1}_{s}$. Thus, $\hat{\theta}^{U}_{s}$ is always less than $\arg\max\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\arg\max\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$, placing it in the first half of both $\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$ The minimum $\hat{\theta}^{U}_{s}>\mu^{0}_{s}+\underline{b}_{M,s}>\arg\max\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$ (second half of $\bm{\Psi}^{0}_{s,(1)}(\theta_{s}))$.
For the optimal non-strategic threshold, $\theta^{U}_{s}<\overline{x}^{0}_{s}<\mu^{0}_{s}+\underline{b}^{0}_{I,s}<\mu^{1}_{s}$, it is always less than $\arg\max\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\arg\max\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$, residing in the first half of $\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$. Its minimum, $\theta^{U}_{s}>\underline{x}^{1}_{s}>\mu^{0}_{s}+\underline{b}_{M,s}$, exceeds $\arg\max\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$, placing it in the second half of $\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$. Since the optimal strategic threshold always exceeds the non-strategic one, and both lie in the first half of $\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$, their impacts can be ordered as $\bm{\Phi}^{y}_{s,(1)}(\theta^{U}_{s})<\bm{\Phi}^{y}_{s,(1)}(\hat{\theta}^{U}_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta^{U}_{s})<\bm{\Psi}^{1}_{s,(1)}(\hat{\theta}^{U}_{s})$. Additionally, as both reside in the second half of $\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$, their impacts order as $\bm{\Psi}^{0}_{s,(1)}(\hat{\theta}^{U}_{s})<\bm{\Psi}^{0}_{s,(1)}(\theta^{U}_{s})$.
Following the same analysis for Type 3 agent responses, the lowest possible optimal unfair strategic threshold, $\hat{\theta}^{U}_{s}>\mu^{0}_{s}+\underline{b}_{M,s}$, exceeds $\arg\max\bm{\Psi}^{0}_{s,(3)}(\theta_{s})$, placing it in the second half of $\bm{\Psi}^{0}_{s,(3)}$. The highest possible $\hat{\theta}^{U}_{s}$, $\overline{x}^{0}_{s}+(\mathbb{T}_{M,s})^{-1}(1-C_{M,s})<\mu^{1}_{s}+\underline{b}_{M,s}$, is less than $\arg\max\bm{\Psi}^{1}_{s,(3)}(\theta_{s})$. Since the non-strategic optimal analysis mirrors the prior section, we can order the impacts as follows: $\bm{\Psi}^{0}_{s,(3)}(\hat{\theta}^{U}_{s})<\bm{\Psi}^{0}_{s,(3)}(\theta^{U}_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta^{U}_{s})<\bm{\Psi}^{1}_{s,(1)}(\hat{\theta}^{U}_{s})$.
A.4 Impacts of relaxing the assumptions of Proposition 8
We start with assumptions $\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}$, and $\overline{x}^{0}_{s}-\underline{x}^{0}_{s}<\mathbf{r}^{y}_{s}-\mathbf{o}^{0}_{s,I}$ imposed within the proposition. If these are violated, for $\alpha_{s}<\xi=\frac{u_{-}}{u_{-}+u_{+}}$ (which, intuitively, means a majority-unqualified population of agents), part (iv) of the proposition would be flipped: $\bm{\Phi}^{0}_{s,(1)}(\theta^{U}_{s})>\bm{\Phi}^{0}_{s,(1)}(\hat{\theta}^{U}_{s})$. This is because under the affordable improvement (Type 1) equilibrium and with a majority-unqualified population of agents, the strategic firm has to increase the threshold considerable to severely restrict successful manipulation by the many unqualified agents, which in turn also reduces the number of unqualified agents, who can afford to opt for improvement compared to the non-strategic firm’s case. This is illustrated in Figure 8, which compares the strategic behavior impact of all agents’ at the optimal policies among different levels of $\alpha_{s}$, the difference from Proposition 8 is shown in the fourth picture from the left.
Figure 8: Impacts of removing $\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}$, and $\overline{x}^{0}_{s}-\underline{x}^{0}_{s}<\mathbf{r}^{y}_{s}-\mathbf{o}^{0}_{s,I}$ from Prop. 8.
Next, if Assumption 3 is relaxed (i.e., $\underline{x}^{1}_{s}<\mu^{0}_{s}<\mu^{1}_{s}<\overline{x}^{0}_{s}$), the conditions of Proposition 8 no longer hold. Specifically, this relaxation always violates $\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{1}_{s}$, and $\underline{x}^{1}_{s}>\mu^{0}_{s}+\underline{b}_{M,s}$ either alone or alongside the other conditions such as $\mu^{0}_{s}+\underline{b}^{0}_{I,s}<\mu^{1}_{s}$ and/or $\overline{x}^{0}_{s}-\underline{x}^{0}_{s}<\mathbf{r}^{y}_{s}-\mathbf{o}^{0}_{s,I}$. For $\alpha_{s}<\xi=\frac{u_{-}}{u_{-}+u_{+}}$ (a majority-unqualified population), part (iv) of Proposition 8 flips, consistent with the prior analysis, yielding $\bm{\Phi}^{0}_{s,(1)}(\theta^{U}s)>\bm{\Phi}^{0}_{s,(1)}(\hat{\theta}^{U}s)$. Here, the strategic firm imposes a high threshold to curb successful manipulation by the many unqualified agents, but this also limits successful manipulation by the few qualified agents. Consequently, part (iii) of the proposition reverses ($\bm{\Psi}^{1}_{s,(1)}(\theta^{U}s)>\bm{\Psi}^{1}_{s,(1)}(\hat{\theta}^{U}_{s})$).
Conversely, when $\alpha_{s}>\xi$ (indicating a majority-qualified population), firms adjust their thresholds to accommodate the many qualified agents, aligning with the reasoning in Proposition 8. The non-strategic firm, unaware of agents’ strategic behavior, lowers its threshold substantially, reducing the need for strategic actions by agents as more are accepted by default. In contrast, the strategic firm, which accounts for this behavior, lowers its threshold less aggressively. This smaller reduction ultimately increases both the number of agents who successfully manipulate to gain acceptance and the number of unqualified agents who can afford to opt for improvement. Accordingly, part (ii) of Proposition 8 flips yielding $\bm{\Psi}^{0}_{s,(1)}(\theta^{U}s)<\bm{\Psi}^{0}_{s,(1)}(\hat{\theta}^{U}_{s})$. This is depicted in Figure 9, which evaluates the impact of strategic behavior across all agents under optimal policies for different levels of $\alpha_{s}$. The deviation from Proposition 8 is evident in the third through fifth pictures from the left.
Figure 9: Impacts of removing Assumption 3 from Proposition 8.
A.5 Type 3 Experiments
Single-group experiment setup.
We model a population of agents from group $s$ with truncated Gaussian feature distributions: $G^{0}_{s}\sim\text{TruncNorm}([0,90],45,15^{2})$ and $G^{1}_{s}\sim\text{TruncNorm}([78,168],123,15^{2})$. The pre-strategic qualification rate $\alpha_{s}\in(0,1)$ varies, with strategic action costs $C_{M,s}=0.2$ and $C_{I,s}=0.8$. Boost distributions are truncated Gaussians: $\tau^{y}_{M,s}(b)\sim\text{TruncNorm}([20,70],45,22)$,
$\tau^{0}_{I,s}(b)\sim\text{TruncNorm}([40,82],61,15)$, and $\tau^{1}_{I,s}(b)\sim\text{TruncNorm}([42,82],62,15)$. We sample $n=1000$ agents based on $\alpha_{s}$, run 50 experiments per $\alpha_{s}$, and report averages.
Strategic vs. non-strategic (unfair) firms.
In Figure 10, the leftmost picture shows strategic optimal thresholds (in the absence of any fairness interventions) exceeding non-strategic ones across all $\alpha_{s}$ levels, aligning with Proposition 8, while the second-left picture confirms higher utility for a strategic firm. The fourth-left picture indicates reduced manipulation by unqualified agents due to anticipating strategic behavior, whereas the third-left picture shows increased manipulation by qualified agents, benefiting the firm, consistent with Proposition 8 parts (ii) and (iii).
Figure 10: Proposition 8 under Type 3 equilibrium
Two-group numerical illustration setup.
We next consider a population consisting of two equal-size groups a and b, with group a being majority-qualified ($\alpha_{a}$ = 0.7) and group b being majority-unqualified ($\alpha_{b}$ = 0.2). The feature distributions for both groups follow truncated Gaussian distributions, with $G^{0}_{a}\sim\text{TruncNorm}([20,110],65,15^{2}),G^{0}_{b}\sim\text{TruncNorm}([0,90],45,15^{2}),G^{1}_{a}\sim\text{TruncNorm}([98,188],143,15^{2}),$ and $G^{1}_{b}\sim\text{TruncNorm}([78,168],123,15^{2})$. We assume both groups have access to the same strategic actions, with $C_{M,s}=0.2$, $C_{I,s}=0.8$, and the same boost distribution as in the "Single-group experiment".
Strategic vs. non-strategic fair firms.
Figure 11 uses contour plots to display firm utility and fairness-constrained utilities (DP and TPR parity) as thresholds $\theta_{a}$ and $\theta_{b}$ vary, marking optimal thresholds (unfair, DP-fair, TPR-fair). Figure 12 shows these thresholds and manipulation effects on utility. The strategic firm, using post-strategic statistics, enforces fairness constraints, unlike the non-strategic firm, which relies on pre-strategic constraints.
As seen in Type 1 equilibrium, both firms raise (resp. lower) thresholds for the majority-qualified (resp. majority-unqualified) group to meet fairness constraints, per Corollary 9 and prior literature, with the strategic firm setting higher thresholds for both groups, consistent with Proposition 8.
Unlike Type 1, comparing DP vs. TPR fairness impacts reveals greater threshold changes under DP for both firms, as seen in Fig. 11. Without improvement actions, $\alpha_{s}$ remains unchanged, forcing firms to lower $\theta_{b}$ and raise $\theta_{a}$ to meet DP constraints, as strategic responses can’t achieve selection parity. For group $b$, the DP-fair strategic firm increases unqualified manipulation more, raising $\bm{\Psi}^{0}_{b,(1)}(\hat{\theta}^{\mathcal{C}}_{b})$ compared to the unfair policy’s high threshold (curb many unqualified manipulation), shown by solid red lines in Figure 12’s first plot. The non-strategic firm lowers $\theta_{b}$ similarly, but with less manipulation increase (Fig. 12’s second plot) due to already high unqualified manipulation. Thus, without improvement actions, fairness requires larger threshold shifts under strategic responses, reducing firm utility.
(a) Nonstrategic utility and fairness constraints
(b) Strategic utility and fairness constraints
Figure 11: Firm’s Utility under agents’ best response Type 3 with the fairness constraint
Figure 12: Strategic behavior’s impact comparison: fair vs unfair policies under Type 3 equilibrium